### v1

# Li Tana,b , Shengrong Wanga,c

Abstract
In this article, we propose a tamed theta Euler-Maruyama (EM) scheme for superlinearly growing neutral McKean-Vlasov stochastic differential delay equations driven by fractional Brownian motion with Hurst exponent $H\in(1/2,1)$. The analytical properties including uniqueness and existence, boundedness of moment and propagation of chaos are investigated. Moreover, the convergence rate of the numerical scheme is established.
Keywords :Neutral McKean-Vlasov SDEs; superlinearly growing; fractional Brownian motion; strong convergence rate; tamed theta EM
1 introduction
Stochastic differential equations (SDEs) with coefficients that depend on the law of the solution were first studied by McKean [1], who was inspired by Kac’s programme in Kinetic Theory [2]. Subsequently, they have become a research hotspot. Bossy and Talay [3, 4] chose to approximate the McKean-Vlasov limit by replicating the behavior with a system of $n$ weakly interacting particles. They also studied the numerical approximation of the solutions to the McKean-Vlasov SDEs. The existence and uniqueness of McKean-Vlasov SDEs were discussed in [5, 6]. Lacker [7] investigated the well-posedness and the propagation of chaos for McKean-Vlasov SDEs. Since the explicit solutions of McKean-Vlasov SDEs are difficult to obtain, many types of McKean-Vlasov SDEs can be approximated by implicit Euler-Maruyama (EM) scheme [8, 16], adaptive EM scheme [9, 11], tamed EM scheme [10, 14], truncated EM scheme [15] and so on. They approximated the solutions of McKean-Vlasov SDEs using numerical methods whose coefficients are globally (or locally) Lipschitz continuous or exhibit super-linear growth.
Recently, there has been a tendency for research to focus on the McKean-Vlasov SDEs driven by fractional Brownian motion. The form is
(1.1) ${\rm d}X_{t}=b(X_{t},\mathcal{L}(X_{t})){\rm d}t+\sigma(\mathcal{L}(X_{t})){\rm d}B_{t}^{H},$
where the coefficients $b:\mathbb{R}^{d}\times\mathcal{P}_{q}(\mathbb{R}^{d})\to\mathbb{R}^{d},\sigma:\mathcal{P}_{q}(\mathbb{R}^{d})\to\mathbb{R}^{d}\otimes\mathbb{R}^{d}$ are Borel-measurable, here, $\mathcal{P}_{q}(\mathbb{R}^{d})$ is the set of probability measures on $\mathbb{R}^{d}$ with finite $q$-th moment. $\mathcal{L}(X_{t})$ is the law of $X$ at $t$ and $B^{H}=\{B_{t}^{H},t\in[0,T]\}$ is a fractional Brownian motion with Hurst exponent $H\in(0,1)$ defined on the probability space $(\Omega,\mathcal{F},\mathbb{P})$. Moreover, $B^{H}$ is a centered Gaussian process with the covariance function
$R_{H}(t,s)=\mathbb{E}(B_{t}^{H}B_{s}^{H})=\frac{1}{2}(t^{2H}+s^{2H}-|t-s|^{2H}),\ \ \forall t,s\in[0,T].$
When $H=1/2$, the fractional Brownian motion $\{B_{t}^{H}\}_{t\geq 0}$ corresponds to a standard Bromnian motion. The self-similarity and long memory properties of fractional Brownian motion make it applicable in various fields, such as atmospheric sciences, financial engineering and others. From then on, scholars conducted studies on (1.1) separately. He et al. [17] established the theory of propagation of chaos and studied the strong convergence of the EM method when the coefficients of (1.1) satisfy the globally Lipschitz condition.
However, few studies considered the case where the coefficients of (1.1) grow super-linearly. Shen [18] demonstrated that, under certain averaging conditions, the solutions of McKean-Vlasov SDEs driven by fractional Brownian motion can be approximated by the solutions of the associated averaged McKean-Vlasov SDEs driven by fractional Brownian motion, in the sense of the mean square convergence, where the coefficients exhibit super-linear growth. Wang et al. [19] and Gao et al. [20] focus on the EM scheme for delayed McKean-Vlasov SDEs driven by fractional Brownian motion, whose drift coefficient contains a polynomial delay term.
Actually, in order to describe the impact of the past, stochastic delay differential equations (SDDEs) and neutral stochastic differential delay equations (NSDDEs) play a significant role in stochastic calculus, describing a wide variety of natural and man-made systems (see, e.g., [21, 22, 23]). The classical EM method is no longer applicable, as the coefficients of SDEs do not exhibit linear growth. It is known that the tamed EM shceme can be used to deal with superlinearly growing coefficients. Deng et al. [24] proposed two types of explicit tamed EM schemes for NSDDEs with superlinearly growing drift and diffusion coefficients and proved strong convergence for each of them, respectively. Cui et al. [25] studied the tamed EM methods to approximate the solutions for a type of of McKean-Vlasov NSDDEs in which the drift coefficients exhibit superlinear growth, and obtained the rate of convergence. We can refer to [23, 27] for more detail about tamed EM scheme in NSDDEs. It is worth noting that Tan and Yuan [28] modified the tamed theta EM scheme and showed the convergence of the numerical solutions, which provides us with research ideas for conducting this paper.
Drawing inspiration from the tamed approach introduced in [28] and previous models of McKean-Vlasov SDEs driven by fractional Brownian motion, we consider a tamed theta EM scheme for MV-NSDDEs driven by fractional Brownian motion with coefficients of superlinear growth, as an extension of (1.1). The form is
(1.2) ${\rm d}(X_{t}-D(X_{t-\tau}))=b(X_{t},X_{t-\tau},\mathcal{L}(X_{t})){\rm d}t+\sigma(\mathcal{L}(X_{t})){\rm d}B_{t}^{H},$
where $\tau>0$ is the delay.
The structure of the rest of the paper is as follows. In Section 2, the mathematical preliminaries on the McKean-Vlasov NSDDEs with superlinear growth driven by fractional Brownian motion are presented. In Section 3, existence and uniqueness of the solution is given. In Section 4, the propagation of chaos for this model is also analyzed. In Section 5, the EM scheme is established and the strong convergence rate is given.
2 Preliminaries
[…]
$\mathcal{P}_{q}(\mathbb{R}^{d})=\bigg{\{}\mu\in\mathcal{P}(\mathbb{R}^{d}):\int_{\mathbb{R}^{d}}|x|^{q}\mu({\rm d}x)<\infty\bigg{\}},$
which is a Polish space under the $L^{q}$-Wasserstein distance [30, Definition 6.1]. Then, for $\mu\in\mathcal{P}_{q}(\mathbb{R}^{d})$, define
$\mathcal{W}_{q}(\mu)=\bigg{(}\int_{\mathbb{R}^{d}}|x|^{q}\mu({\rm d}x)\bigg{)}^{1/q}$
For $\mu,\nu\in\mathcal{P}_{q}(\mathbb{R}^{d})$, we define the Wasserstein distance on $\mathcal{P}_{q}(\mathbb{R}^{d})$ by
$\mathbb{W}_{q}(\mu,\nu)=\bigg{(}\inf_{\pi\in\Pi(\mu,\nu)}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}|x-y|^{q}\pi({\rm d}x,{\rm d}y)\bigg{)}^{1/q},$
where $\Pi(\mu,\nu)$ is the set of probability measures on $\mathbb{R}^{d}\times\mathbb{R}^{d}$ with marginals $\mu$ and $\nu$. In other words, $\pi\in\Pi(\mu,\nu)$ is a probability measure on $\mathbb{R}^{d}\times\mathbb{R}^{d}$ such that $\pi(\cdot\times\mathbb{R}^{d})=\mu(\cdot)$ and $\pi(\cdot\times\mathbb{R}^{d})=\nu(\cdot)$. For $11/2<H<$ 1, we put
(2.1) $\phi(s,t)=\phi_{H}(s,t)=H(2H-1)|s-t|^{2H-2},\ \ s,t\in\mathbb{R},$
and recall that for $s,t>0$,
$\int_{0}^{t}\int_{0}^{s}\phi(u,v){\rm d}u{\rm d}v=R_{H}(t,s).$
Let $\mathcal{S}(\mathbb{R})$ be the Schwartz space of rapidly decreasing smooth functions on $R$, and if $b\in\mathcal{S}(\mathbb{R})$, denote
$\parallel f\parallel_{H}^{2}:=\int_{\mathbb{R}}\int_{\mathbb{R}}f(s)f(t)\phi(s,t){\rm d}s{\rm d}t<\infty.$
If we equip $\mathcal{S}(\mathbb{R})$ with the inner product
$\langle f,g\rangle_{H}:=\int_{\mathbb{R}}\int_{\mathbb{R}}f(s)g(t)\phi(s,t){\rm d}s{\rm d}t,\ \ f,g\in\mathcal{S}(\mathbb{R}),$
then the completion of $\mathcal{S}(\mathbb{R})$, denoted by $L_{\phi}^{2}(\mathbb{R})$,
$L_{\phi}^{2}(\mathbb{R}_{+})=\bigg{\{}g|g:\mathbb{R}_{+}\to\mathbb{R},\ \ \ |g|_{\phi}^{2}:=\int_{0}^{\infty}\int_{0}^{\infty}g(s)g(t)\phi(s,t){\rm d}s{\rm d}t<\infty\bigg{\}}.$
Definition 2.1.
[32, Definition 3.5.1] Let $g\in L_{\phi}^{2}(\mathbb{R})$. The $\phi$-derivative of a stochastic variable $F\in L^{p}$ in the direction of $\Phi_{g}$ is defined by
$D_{\Phi_{g}}F(\omega)=\lim_{\delta\to 0}\frac{1}{\delta}\bigg{\{}F\bigg{(}\omega+\delta\int_{0}^{\cdot}(\Phi_{g})(u){\rm d}u\bigg{)}-F(\omega)\bigg{\}},$
if the limit exists in $L^{p}$. Furthermore, if there is a process $(D_{s}^{\phi}F,s\geq 0)$ such that
$D_{\Phi_{g}}F=\int_{0}^{\infty}D_{s}^{\phi}Fg_{s}{\rm d}s\ \ \ {\rm almost\ surely},$
for all $g\in L_{\phi}^{2}(\mathbb{R})$, then $F$ is said to be $\phi$-differentiable.
For SDEs driven by fractional Brownian motion
(2.2) ${\rm d}X_{t}=G(t,X_{t}){\rm d}t+F(t,X_{t}){\rm d}B_{t}^{H}.$
Let $C^{2,1}(\mathbb{R}\times\mathbb{R}_{+};\mathbb{R})$ be the family of all real-valued functions $V(x,t)$ defined on $\mathbb{R}\times\mathbb{R}_{+}$, which is twice continuously differentiable in $x$ and once in $t$. Referring to the Itô formula for fractional Brownian motion given by [32, 33],
${\rm d}V(t,X_{t})=$ $\frac{\partial V(t,X_{t})}{\partial t}{\rm d}t+\frac{\partial V(t,X_{t})}{\partial X}G(t,X_{t}){\rm d}t+\frac{\partial V(t,X_{t})}{\partial X}F(t,X_{t}){\rm d}B_{t}^{H}$
$+\frac{\partial^{2}V(t,X_{t})}{\partial X^{2}}F(t,X_{t})D_{t}^{\phi}X_{t}{\rm d}t,$
or formally,
${\rm d}V(t,X_{t})=$ $\frac{\partial V(t,X_{t})}{\partial t}{\rm d}t+\frac{\partial V(t,X_{t})}{\partial X}G_{t,X_{t}}{\rm d}t+\frac{\partial V(t,X_{t})}{\partial X}F(t,X_{t}){\rm d}B_{t}^{H}$
$+\frac{\partial^{2}V(t,X_{t})}{\partial X^{2}}F(t,X_{t})\int_{0}^{t}\phi(t,r)F(r,X_{r}){\rm d}r{\rm d}t.$
In addition, [34], [35] also discussed the fractional Itô formula. Now, we are going to give the fractional Itô formula as follows.
Remark 2.1.
(Fractional Itô formula) Assume that the stochastic process $X_{t}$ follows (2.2) with $11/2<H<$ 1. We have
$V(t,X_{t})$
$=$ $V(X_{0})+\int_{0}^{t}\bigg{[}\frac{\partial V(s,X_{s})}{\partial t}+\frac{\partial V(s,X_{s})}{\partial X}G(s,X_{s})$
$+H(2H-1)\frac{\partial^{2}V(s,X_{s})}{\partial X^{2}}\int_{0}^{s}(s-r)^{2H-2}F^{2}(r,X_{r}){\rm d}r\bigg{]}{\rm d}s+\int_{0}^{t}\frac{\partial V(s,X_{s})}{\partial X}F(s,X_{s}){\rm d}B_{s}^{H}$
$=$ $V(X_{0})+\int_{0}^{t}\bigg{[}\frac{\partial V(s,X_{s})}{\partial t}+\frac{\partial V(s,X_{s})}{\partial X}G(s,X_{s})+H\frac{\partial^{2}V(s,X_{s})}{\partial X^{2}}s^{2H-1}F^{2}(s,X_{s})\bigg{]}{\rm d}s$
$+\int_{0}^{t}\frac{\partial V(s,X_{s})}{\partial X}F(s,X_{s}){\rm d}B_{s}^{H}.$
Since the kernel function $\phi$ is defined only for $11/2<H<$ 1, otherwise, $\phi$ has no meaning. We assume that $11/2<H<$ 1 throughout this paper. In this paper, we consider the following $d$-dimensional McKean-Vlasov NSDDEs driven by fractional Brownian motion as follows:
(2.5) $\left\{\begin{array}[]{ll}{\rm d}(X_{t}-D(X_{t-\tau}))=b(X_{t},X_{t-\tau},\mathcal{L}_{X_{t}}){\rm d}t+\sigma(\mathcal{L}_{X_{t}}){\rm d}B_{t}^{H},\ \ t>0,\\ X_{t}=\xi=\{\xi(t):-\tau\leq t\leq 0\}\in L_{\mathcal{F}_{0}}^{p}([-\tau,0];\mathbb{R}^{d}),\ p\geq\theta\geq 1\end{array}\right.$
where $b:\mathbb{R}^{d}\times\mathbb{R}^{d}\times\mathcal{P}_{q}(\mathbb{R}^{d})\to\mathbb{R}^{d}$, $\sigma:\mathcal{P}_{q}(\mathbb{R}^{d})\to\mathbb{R}^{d}\otimes\mathbb{R}^{d}$, $\mathcal{L}_{X_{t}}$ denotes the law of $X_{t}$. Throughout the paper, we shall denote $C$ by a generic positive constant, whose value may change from line to line. For any $x,y,\overline{x},\overline{y}$, we impose the following assumptions.
Assumption 2.2.
For any $s,t\in[-\tau,0]$ and $p>0$, there is a positive constant $K_{0}$ such that
$\mathbb{E}\bigg{(}\sup_{-\tau\leq s,t\leq 0}|\xi(s)-\xi(t)|^{p}\bigg{)}\leq K_{0}|s-t|^{p}.$
Assumption 2.3.
$D(0)=0$, and there exist positive constants $\lambda<1$ such that
$|D(x)-D(\overline{x})|\leq\lambda|x-\overline{x}|.$
Assumption 2.4.
There exist positive constants $l\geq 1$ and $K_{2},K_{3}$ such that
$\big{\langle}x-D(y)-\overline{x}+D(\overline{y}),b(x,y,\mu)-b(\overline{x},\overline{y},\nu)\big{\rangle}\leq K_{2}(|x-\overline{x}|^{2}+|y-\overline{y}|^{2}+(\mathbb{W}_{q}(\mu,\nu))^{2}),$
and
$|b(x,y,\mu)-b(\overline{x},\overline{y},\nu)|\leq K_{3}[(1+|x|^{l}+|\overline{x}|^{l}+|y|^{l}+|\overline{y}|^{l})(|x-\overline{x}|+|y-\overline{y}|)+\mathbb{W}_{q}(\mu,\nu)],$
for $x,y,\overline{x},\overline{y}\in\mathbb{R}^{d}$, $\mu,\nu\in\mathcal{P}_{q}(\mathbb{R}^{d})$.
Assumption 2.5.
There exist positive constants $K_{4},K_{5}$ such that
$\parallel\sigma(\mu)-\sigma(\nu)\parallel\leq K_{4}\mathbb{W}_{q}(\mu,\nu),$
and,
$|b(0,0,\mu)|\vee\parallel\sigma(\mu)\parallel\leq K_{5}(1+\mathbb{W}_{q}(\mu,\delta_{0})),$
for $\mu,\nu\in\mathcal{P}_{q}(\mathbb{R}^{d})$. Then, for initial experience distribution $\delta_{0}$
$|b(0,0,\delta_{0})|\vee\parallel\sigma(\delta_{0})\parallel\leq K_{6}.$
Remark 2.6.
Let Assumption 2.4 hold, then, for any $x,y\in\mathbb{R}^{d}$, we have
$|b(x,y,\mu)|\leq$ $|b(x,y,\mu)-b(0,0,\mu)|+|b(0,0,\mu)|$
$\leq$ $K_{3}(1+|x|^{l}+|y|^{l})(|x|+|y|)+K_{5}(1+\mathbb{W}_{q}(\mu,\delta_{0}))$
$\leq$ $C(1+|x|^{l+1}+|y|^{l+1}+\mathbb{W}_{q}(\mu,\delta_{0})),$
where $C=K_{3}\vee K_{5}$. Moreover, by Young’s inequality, we get
$(x-D(y))b(x,y,\mu)=$ $(x-D(y)-0+D(0))(b(x,y,\mu)-b(0,0,\delta_{0})+b(0,0,\delta_{0}))$
$\leq$ $K_{2}(|x|^{2}+|y|^{2}+\mathbb{W}_{q}^{2}(\mu,\delta_{0}))+(|x|+D(y))|b(0,0,\delta_{0})|$
$\leq$ $C(1+|x|^{2}+|y|^{2}+\mathbb{W}_{q}^{2}(\mu,\delta_{0})),$
where $C=K_{2}\vee(K_{2}+1)\vee(K_{2}+\lambda)\vee(K_{2}+K_{5}^{2})$.
Lemma 2.7.
[31, Lemma 4.1] Let $p>1$, $\varepsilon>0$ and $x,y\in\mathbb{R}$. Then
$|x+y|^{p}\leq\bigg{(}1+\varepsilon^{\frac{1}{p-1}}\bigg{)}^{p-1}\bigg{(}|x|^{p}+\frac{|y|^{p}}{\varepsilon}\bigg{)}.$
Especially for $\varepsilon=((1-a)/a)^{p-1}$, where $p\geq 2$ and $10<a<$ 1, we get
$|x+y|^{p}\leq\frac{1}{a^{p-1}}|x|^{p}+\frac{1}{(1-a)^{p-1}}|y|^{p}.$
3 Existence and Uniqueness
Since we are studying equations with super-linearly growing coefficients, it is necessary to consider the Carathéodory approximation. Now we define the Carathéodory approximation of (2.5) as follows. For every integer $n\geq 2/\tau$, define $X_{t}^{n}$ on $[-\tau,T]$ by
$X_{t}^{n}=\xi(t),\ \ t\in[-\tau,0],$
and
(3.1) $X_{t}^{n}-D(X_{t-\tau}^{n})=\xi(0)-D(\xi(-\tau))+\int_{0}^{t}b\big{(}X_{s-1/n}^{n},X_{s-\tau}^{n},\mathcal{L}_{X_{s-1/n}^{n}}\big{)}{\rm d}s+\int_{0}^{t}\sigma\big{(}\mathcal{L}_{X_{s-1/n}^{n}}\big{)}{\rm d}B_{s}^{H},$
for $t\in[0,T]$. It is important to note that each $X_{t}^{n}$ can be determined explicitly by the stepwise iterated Itô integrals over the intervals $[0,1/n]$, $(1/n,2/n]$ etc. To get the final results, we give the following lemmas.
Lemma 3.1.
Let Assumptions 2.2-2.5 hold. Then for $\overline{p}>1/H$ and $n\geq 2/\tau$, we get
$\mathbb{E}\bigg{(}\sup_{0\leq t\leq T}|X_{t}^{n}|^{\overline{p}}\bigg{)}\leq C.$
Proof.
By Assumption 2.4 and Lemma 2.7, we get
$|X_{t}^{n}|^{\overline{p}}=$ $|D(X_{t-\tau}^{n})+X_{t}^{n}-D(X_{t-\tau}^{n})|^{\overline{p}}$
$\leq$ $\frac{1}{\lambda^{p-1}}|D(X_{t-\tau}^{n})|^{\overline{p}}+\frac{1}{(1-\lambda)^{{\overline{p}}-1}}|X_{t}^{n}-D(X_{t-\tau}^{n})|^{\overline{p}}$
$\leq$ $\lambda|X_{t-\tau}^{n}|^{\overline{p}}+\frac{1}{(1-\lambda)^{p-1}}|X_{t}^{n}-D(X_{t-\tau}^{n})|^{\overline{p}},$
for any $t\in[0,T]$. Then
$\sup_{0\leq r\leq t}|X_{r}^{n}|^{\overline{p}}\leq$ $\lambda\parallel\xi\parallel^{p}+\lambda\sup_{0\leq r\leq t}|X_{r}^{n}|^{\overline{p}}+\frac{1}{(1-\lambda)^{{\overline{p}}-1}}\sup_{0\leq r\leq t}|X_{r}^{n}-D(X_{r-\tau}^{n})|^{\overline{p}},$
we may compute that
(3.2) $\sup_{0\leq r\leq t}|X_{r}^{n}|^{\overline{p}}\leq\frac{\lambda}{1-\lambda}\parallel\xi\parallel^{\overline{p}}+\frac{1}{(1-\lambda)^{\overline{p}}}\sup_{0\leq r\leq t}|X_{r}^{n}-D(X_{r-\tau}^{n})|^{\overline{p}}$
By Assumption 2.3, we get
$\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|X_{s}^{n}|^{\overline{p}}\bigg{)}$
$\leq$ $3^{\overline{p}-1}\mathbb{E}|\xi(0)-D(\xi(-\tau))|^{\overline{p}}+3^{\overline{p}-1}\mathbb{E}\bigg{(}\sup_{s\in[0,t]}\bigg{\lvert}\int_{0}^{s}b\left(X_{r-1/n}^{n},X_{r-\tau}^{n},\mathcal{L}_{X_{r-1/n}^{n}}\right){\rm d}r\bigg{\rvert}^{\overline{p}}\bigg{)}$
$+3^{\overline{p}-1}\mathbb{E}\bigg{(}\sup_{s\in[0,t]}\bigg{\lvert}\int_{0}^{s}\sigma\left(\mathcal{L}_{X_{r-1/n}^{n}}\right){\rm d}B_{r}^{H}\bigg{\rvert}^{\overline{p}}\bigg{)}$
$\leq$ $C+C\mathbb{E}\int_{0}^{t}\left[1+|X_{r-1/n}^{n}|^{(l+1)\overline{p}}+|X_{r-\tau}^{n}|^{(l+1){\overline{p}}}+\left(\mathbb{W}_{\theta}(\mathcal{L}_{X_{r-1/n}^{n}},\delta_{0})\right)^{\overline{p}}\right]{\rm d}r$
$+C\mathbb{E}\bigg{(}\sup_{s\in[0,t]}\bigg{\lvert}\int_{0}^{s}\sigma(\mathcal{L}_{X_{r-1/n}^{n}}){\rm d}B_{r}^{H}\bigg{\rvert}^{\overline{p}}\bigg{)}.$
Using the techniques of [12, Theorem 3.1]. Denote $\int_{s}^{t}(t-\beta)^{-\beta}(r-s)^{\beta-1}{\rm d}r:=C(\beta)$. Since for $\overline{p}H>1$, we can take some $\beta$ such that $1-H<\beta<1-1/{H}$, then by the stochastic Fubini theorem, the Hölder inequality, and combining with Lemma 2.4, we get
(3.3) $\mathbb{E}\bigg{(}\sup_{s\in[0,t]}\bigg{\lvert}\int_{0}^{s}\sigma(\mathcal{L}_{X_{r-1/n}^{n}}){\rm d}B_{r}^{H}\bigg{\rvert}^{\overline{p}}\bigg{)}$
$\leq$ $\frac{C(\beta)^{-{\overline{p}}}}{({\overline{p}}-1-\beta{\overline{p}})^{{\overline{p}}-1}}t^{{\overline{p}}-1-\beta{\overline{p}}}\int_{0}^{t}\mathbb{E}\bigg{|}\int_{0}^{s}(s-r)^{\beta-1}\sigma(\mathcal{L}_{X_{r-1/n}^{n}}){\rm d}B_{r}^{H}\bigg{|}^{{\overline{p}}}{\rm d}s$
$\leq$ $Ct^{{\overline{p}}-1-\beta{\overline{p}}}\int_{0}^{t}\left(\int_{0}^{s}(s-r)^{\frac{\beta-1}{H}}\parallel\sigma(\mathcal{L}_{X_{r-1/n}^{n}})\parallel^{\frac{1}{H}}{\rm d}r\right)^{\overline{p}H}{\rm d}s$
$\leq$ $Ct^{\overline{p}H-1}\int_{0}^{t}\parallel\sigma(\mathcal{L}_{X_{r-1/n}^{n}})\parallel^{\overline{p}}{\rm d}r$
$\leq$ $Ct^{\overline{p}H-1}L^{\overline{p}}\int_{0}^{t}\left[1+\mathbb{W}_{\theta}(\mathcal{L}_{X_{r-1/n}^{n}},\delta_{0})\right]^{\overline{p}}{\rm d}r.$
Referring to the proof in [6, Proposition 3.4], for the $q$-Wasserstein metric, we get
$\left(\mathbb{W}_{q}(\mathcal{L}_{X_{r-1/n}^{n}},\delta_{0})\right)^{\overline{p}}=\left(\mathcal{W}_{q}(\mathcal{L}_{X_{r-1/n}^{n}})\right)^{\overline{p}}\leq\mathbb{E}|X_{r-1/n}^{n}|^{\overline{p}}.$
To sum up, we obtain
$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}|X_{t}^{n}-D(X_{t-\tau}^{n})|^{\overline{p}}\bigg{)}\leq$ $C+C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{0\leq u\leq s}|X_{u-\tau}^{n}|^{(l+1)\overline{p}}\bigg{)}{\rm d}s,$
together with (3.2), we get
$\mathbb{E}\bigg{(}\sup_{0\leq r\leq t}|X_{r}^{n}|^{\overline{p}}\bigg{)}\leq$ $C+C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{0\leq u\leq s}|X_{u-\tau}^{n}|^{(l+1)\overline{p}}\bigg{)}{\rm d}s.$
We set up a finite sequence $\{u_{1},u_{2},\cdots,u_{[T/\tau]+1}\}$ where $u_{i}=([T/\tau]+2-i)\overline{p}(l+1)^{[T/\tau]+1-i}$, and $[T/\tau]$ denotes the integer part of $T/\tau$. It is easy to see that $u_{i}\geq\overline{p}$, $u_{[T/\tau]+1}=\overline{p}$ and $(l+1)u_{i+1}<u_{i}$ for $i=1,2,\cdots,[T/\tau]$. For $s\in[0,\tau]$, we immediately get
$\mathbb{E}\bigg{(}\sup_{0\leq r\leq t}|X_{r}^{n}|^{\overline{p}}\bigg{)}\leq C,$
since $\xi\in L_{\mathcal{F}_{0}}^{\overline{p}}([-\tau,0];\mathbb{R}^{d})$. For $s\in[0,2\tau]$, by the Hölder inequality, we get
$\mathbb{E}\bigg{(}\underset{s\in[0,2\tau]}{\sup}|X_{s}^{n}|^{u_{2}}\bigg{)}\leq C+C\int_{0}^{2\tau}\left[\mathbb{E}\bigg{(}\sup_{s\in[0,r]}|X_{s-\tau}^{n}|^{u_{1}}\bigg{)}\right]^{\frac{(l+1)u_{2}}{u_{1}}}{\rm d}r\leq C.$
By induction, we can derive the assertion. ∎
Lemma 3.2.
Let Assumptions 2.2-2.5 hold. Then for $0\leq s\leq t\leq T$ with $t-s<1$, and $p>1/H$, $(l+1)p\leq\overline{p}$, $n\geq 2/\tau$, we have
$\mathbb{E}|X_{t}^{n}-X_{s}^{n}|^{p}\leq C[(t-s)^{p}+(t-s)^{pH}],$
where $C$ is a positive constant depending on $p,H,\xi,T,L$ but independent of $s$ and $t$.
Proof.
By (3.1), we may compute
$X_{t}^{n}-X_{s}^{n}=$ $D(X_{t-\tau}^{n})-D(X_{s-\tau}^{n})+\int_{s}^{t}b\left(X^{n}_{r-1/n},X^{n}_{r-\tau},\mathcal{L}_{X_{r-1/n}^{n}}\right){\rm d}r+\int_{s}^{t}\sigma(\mathcal{L}_{X_{r-1/n}^{n}}){\rm d}B_{r}^{H}.$
By the elementary inequality, the Hölder inequality, Assumptions 2.3, 2.4, 2.5 and using same techniques as (3.3), we get
$\mathbb{E}|X_{t}^{n}-X_{s}^{n}|^{p}$
$\leq$ $3^{p-1}\mathbb{E}|D(X_{t-\tau}^{n})-D(X_{s-\tau}^{n})|^{p}+3^{p-1}\mathbb{E}\bigg{\lvert}\int_{s}^{t}b(X_{r-1/n}^{n},X_{r-\tau}^{n},\mathcal{L}_{X_{r-1/n}^{n}}){\rm d}r\bigg{\rvert}^{p}$
$+3^{p-1}\mathbb{E}\bigg{\lvert}\int_{s}^{t}\sigma(\mathcal{L}_{X_{r-1/n}^{n}}){\rm d}B_{r}^{H}\bigg{\rvert}^{p}$
$\leq$ $3^{p-1}\bigg{[}\lambda^{p}\mathbb{E}|X_{t-\tau}^{n}-X_{s-\tau}^{n}|^{p}+(t-s)^{p-1}$
×$\times\mathbb{E}\int_{s}^{t}C^{p}\bigg{(}1+|X_{r-1/n}^{n}|^{l+1}+|X_{r-\tau}^{n}|^{l+1}+\mathbb{W}_{q}(\mathcal{L}_{X_{r-1/n}^{n}},\delta_{0})\bigg{)}^{p}\bigg{]}{\rm d}r$
$+C(t-s)^{pH-1}\mathbb{E}\int_{s}^{t}\bigg{(}1+\mathbb{W}_{q}(\mathcal{L}_{X_{r-1/n}^{n}},\delta_{0})\bigg{)}^{p}\bigg{]}.$
For $0\leq s\leq t\leq\tau$, by Assumption 2.2, we get
$\mathbb{E}|X_{t}^{n}-X_{s}^{n}|^{p}$
$\leq$ $3^{p-1}\lambda^{p}\mathbb{E}|\xi(t-\tau)-\xi(s-\tau)|^{p}+C(t-s)^{p-1}\int_{s}^{t}\mathbb{E}\bigg{(}\sup_{u\in[-\tau,r]}|X_{u}^{n}|^{(l+1)p}\bigg{)}{\rm d}r$
$+C(t-s)^{pH-1}\int_{s}^{t}\mathbb{E}\bigg{(}\sup_{u\in[-\tau,r]}|X_{u}^{n}|^{p}\bigg{)}{\rm d}r$
$\leq$ $3^{p-1}\lambda^{p}K_{0}|t-s|^{p}+C(t-s)^{p-1}\int_{s}^{t}\mathbb{E}\bigg{(}\sup_{u\in[-\tau,r]}|X_{u}^{n}|^{(l+1)p}\bigg{)}{\rm d}r$
$+C(t-s)^{pH-1}\int_{s}^{t}\mathbb{E}\bigg{(}\sup_{u\in[-\tau,r]}|X_{u}^{n}|^{p}\bigg{)}{\rm d}r.$
Since $(l+1)p\leq\overline{p}$, we can applying Lemma 3.1 to obtain the assertion. For $\tau\leq s\leq t\leq T$, noting that $01-3^{p-1}\lambda^{p}>$ 0, we get the desired assertion by applying Lemma 3.1 again. ∎
Theorem 3.3.
Let Assumption 2.4 hold. Then for $p>1/H$ and $(l+1)p\leq\overline{p}$, (2.5) admits a global unique solution.
Proof.
We split the proof into two steps.
Step 1. Existence. For $m>n\geq 2/\tau$ and $t\in[0,T]$, by (2.5), we may compute
$\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|X^{m}_{s}-X_{s}^{n}|^{p}\bigg{)}$
$\leq$ $3^{p-1}\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|D(X_{s-\tau}^{m})-D(X_{s-\tau}^{n})|^{p}\bigg{)}$
$+3^{p-1}\mathbb{E}\bigg{(}\sup_{s\in[0,t]}\bigg{\lvert}\int_{0}^{s}\bigg{[}b\big{(}X_{r-1/m}^{m},X_{r-\tau}^{m},\mathcal{L}_{X_{r-1/m}^{m}}\big{)}-b\big{(}X_{r-1/n}^{n},X_{r-\tau}^{n},\mathcal{L}_{X_{r-1/n}^{n}}\big{)}\bigg{]}{\rm d}r\bigg{\rvert}^{p}\bigg{)}$
$+3^{p-1}\mathbb{E}\bigg{(}\sup_{s\in[0,t]}\bigg{\lvert}\int_{0}^{s}\bigg{[}\sigma\big{(}\mathcal{L}_{X_{r-1/m}^{m}}\big{)}-\sigma\big{(}\mathcal{L}_{X_{r-1/n}^{n}}\big{)}\bigg{]}{\rm d}B_{r}^{H}\bigg{\rvert}^{p}\bigg{)}$
=:absent:=:= : $W_{1}+W_{2}+W_{3},$
By Assumption 2.3, we get
$\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|D(X_{s-\tau}^{m})-D(X_{s-\tau}^{n})|^{p}\bigg{)}\leq\lambda^{p}\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|X_{s-\tau}^{m}-X_{s-\tau}^{n}|^{p}\bigg{)}.$
By the Hölder inequality, Assumption 2.4, we get
$W_{2}=$ $3^{p-1}\mathbb{E}\bigg{(}\sup_{s\in[0,t]}\bigg{\lvert}\int_{0}^{s}\bigg{[}b\big{(}X_{r-1/m}^{m},X_{r-\tau}^{m},\mathcal{L}_{X_{r-1/m}^{m}}\big{)}-b\big{(}X_{r-1/n}^{n},X_{r-\tau}^{n},\mathcal{L}_{X_{r-1/n}^{n}}\big{)}\bigg{]}{\rm d}r\bigg{\rvert}^{p}\bigg{)}$
$\leq$ $(3t)^{p-1}\mathbb{E}\int_{0}^{t}\bigg{\lvert}b\big{(}X_{r-1/m}^{m},X_{r-\tau}^{m},\mathcal{L}_{X_{r-1/m}^{m}}\big{)}-b\big{(}X_{r-1/n}^{n},X_{r-\tau}^{n},\mathcal{L}_{X_{r-1/n}^{n}}\big{)}\bigg{\rvert}^{p}{\rm d}r$
$\leq$ $(3t)^{p-1}K_{3}\mathbb{E}\int_{0}^{t}\bigg{[}\bigg{(}1+|X_{r-1/m}^{m}|^{l}+|X_{r-1/n}^{n}|^{l}+|X_{r-\tau}^{m}|^{l}+|X_{r-\tau}^{n}|^{l}\bigg{)}$
×$\times\bigg{(}|X_{r-1/m}^{m}-X_{r-1/n}^{n}|+|X_{r-\tau}^{m}-X_{s-\tau}^{n}|\bigg{)}+\mathbb{W}_{q}\big{(}\mathcal{L}_{X_{r-1/m}^{m}},\mathcal{L}_{X_{r-1/n}^{n}}\big{)}\bigg{]}^{p}{\rm d}r$
$\leq$ $C\mathbb{E}\int_{0}^{t}\bigg{(}1+|X_{r-1/m}^{m}|^{l}+|X_{r-1/n}^{n}|^{l}+|X_{r-\tau}^{m}|^{l}+|X_{r-\tau}^{n}|^{l}\bigg{)}^{p}$
×$\times\bigg{(}|X_{r-1/m}^{m}-X_{r-1/n}^{n}|+|X_{r-\tau}^{m}-X_{s-\tau}^{n}|\bigg{)}^{p}{\rm d}r$
$+C\mathbb{E}\int_{0}^{t}|X_{r-1/m}^{m}-X_{r-1/n}^{n}|^{p}{\rm d}r.$
Since for $l>1$, we have $2pl\leq(l+1)p\leq\overline{p}$, then by the Hölder inequality, Lemmas 3.1, 3.2, we get
$W_{2}\leq$ $C\int_{0}^{t}\bigg{[}\mathbb{E}\bigg{(}1+|X_{r-1/m}^{m}|^{l}+|X_{r-1/n}^{n}|^{l}+|X_{r-\tau}^{m}|^{l}+|X_{r-\tau}^{n}|^{l}\bigg{)}^{2p}\bigg{]}^{1/2}$
×$\times\bigg{[}\mathbb{E}\bigg{(}|X_{r-1/m}^{m}-X_{r-1/n}^{n}|+|X_{r-\tau}^{m}-X_{s-\tau}^{n}|\bigg{)}^{2p}\bigg{]}^{1/2}{\rm d}r+C\mathbb{E}\int_{0}^{t}|X_{r-1/m}^{m}-X_{r-1/n}^{n}|^{p}{\rm d}r$
$\leq$ $C\int_{0}^{t}\bigg{[}\mathbb{E}\bigg{(}|X_{r-1/m}^{m}-X_{r-1/n}^{n}|+|X_{r-\tau}^{m}-X_{r-\tau}^{n}|\bigg{)}^{2p}\bigg{]}^{1/2}{\rm d}r$
$+C\bigg{[}\bigg{(}\frac{1}{n}-\frac{1}{m}\bigg{)}^{p}+\bigg{(}\frac{1}{n}-\frac{1}{m}\bigg{)}^{pH}\bigg{]}.$
By Assumption 2.5, Lemmas 3.1, 3.2 and the same techniques as (3.3), we get
$W_{3}=$ $3^{p-1}\mathbb{E}\bigg{[}\sup_{s\in[0,t]}\bigg{\lvert}\int_{0}^{s}\bigg{(}\sigma(\mathcal{L}_{X_{r-1/m}^{m}})-\sigma(\mathcal{L}_{X_{r-1/n}^{n}})\bigg{)}{\rm d}B_{r}^{H}\bigg{\rvert}^{p}\bigg{]}$
$\leq$ $C\mathbb{E}\int_{0}^{t}|\sigma(\mathcal{L}_{X_{r-1/m}^{m}})-\sigma(\mathcal{L}_{X_{r-1/n}^{n}})|^{p}{\rm d}r$
$\leq$ $C\mathbb{E}\int_{0}^{t}K_{4}^{p}(\mathbb{W}_{q}(\mathcal{L}_{X_{r-1/m}^{m}}-\mathcal{L}_{X_{r-1/n}^{n}}))^{p}{\rm d}r$
$\leq$ $C\int_{0}^{t}\mathbb{E}|X_{r-1/m}^{m}-X_{r-1/n}^{n}|^{p}{\rm d}r$
$\leq$ $C\int_{0}^{t}\mathbb{E}|X_{r-1/m}^{m}-X_{r-1/n}^{m}|^{p}{\rm d}r+C\int_{0}^{t}\mathbb{E}|X_{r-1/n}^{m}-X_{r-1/n}^{n}|^{p}{\rm d}r$
$\leq$ $C\bigg{[}\bigg{(}\frac{1}{n}-\frac{1}{m}\bigg{)}^{p}+\bigg{(}\frac{1}{n}-\frac{1}{m}\bigg{)}^{pH}\bigg{]}+C\int_{0}^{t}\mathbb{E}|X_{r-1/n}^{m}-X_{r-1/n}^{n}|^{p}{\rm d}r.$
Combining $W_{1},W_{2},W_{3}$, we get
$\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|X^{m}_{s}-X_{s}^{n}|^{p}\bigg{)}$
$\leq$ $\lambda^{p}\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|X_{s-\tau}^{m}-X_{s-\tau}^{n}|^{p}\bigg{)}+C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{u\in[0,r]}|X_{u}^{m}-X_{u}^{n}|^{p}\bigg{)}{\rm d}r$
$+C\bigg{[}\bigg{(}\frac{1}{n}-\frac{1}{m}\bigg{)}^{p}+\bigg{(}\frac{1}{n}-\frac{1}{m}\bigg{)}^{pH}\bigg{]}+C\int_{0}^{t}\bigg{[}\mathbb{E}\bigg{(}\sup_{u\in[0,r]}|X_{u-\tau}^{m}-X_{u-\tau}^{n}|\bigg{)}^{2p}\bigg{]}^{1/2}{\rm d}r.$
Noting that $\lambda<1$, by the Gronwall inequality, we get
$\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|X^{m}_{s}-X_{s}^{n}|^{p}\bigg{)}$
$\leq$ $C\bigg{[}\bigg{(}\frac{1}{n}-\frac{1}{m}\bigg{)}^{p}+\bigg{(}\frac{1}{n}-\frac{1}{m}\bigg{)}^{pH}\bigg{]}+C\int_{0}^{t}\bigg{[}\mathbb{E}\bigg{(}\sup_{u\in[0,r]}|X_{u-\tau}^{m}-X_{u-\tau}^{n}|\bigg{)}^{2p}\bigg{]}^{1/2}{\rm d}r.$
Define a sequence $\{u_{1},u_{2},\cdots,u_{[T/\tau]+1}\}$ where $u_{i}=([T/\tau]+2-i)p2^{[T/\tau]+1-i}$, and $[T/\tau]$ denotes the integer part of $T/\tau$. It is easy to see that $u_{i}\geq p$, $u_{[T/\tau]+1}=p$ and $2u_{i+1}<u_{i}$ for $i=1,2,\cdots,[T/\tau]$. For $t\in[0,\tau]$, we get
(3.4) $\mathbb{E}\bigg{(}\sup_{s\in[0,\tau]}|X^{m}_{s}-X_{s}^{n}|^{u_{1}}\bigg{)}\to 0,\ {\rm for}\ m,n\to\infty.$
For $t\in[0,2\tau]$, by the Hölder inequality, we get
$\mathbb{E}\bigg{(}\underset{s\in[0,2\tau]}{\sup}|X^{m}_{s}-X_{s}^{n}|^{u_{2}}\bigg{)}$
[…]
by (3.4),
$\mathbb{E}\bigg{(}\underset{s\in[0,2\tau]}{\sup}|X^{m}_{s}-X_{s}^{n}|^{u_{2}}\bigg{)}\to 0,\ {\rm for}\ m,n\to\infty.$
By induction, we can see that for $t\in[0,T]$,
(3.5) $\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|X^{m}_{s}-X_{s}^{n}|^{p}\bigg{)}\to 0,\ {\rm for}\ m,n\to\infty.$
Consequently, $(X^{n})_{n\geq 2/\tau}$ is a Cauchy sequence for $p>1/H$. Denote the limit of $(X^{n})$ by $X$ and letting $m\to\infty$ in (3.5),
(3.6) $\lim_{n\to\infty}\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|X_{s}-X_{s}^{n}|^{p}\bigg{)}=0.$
Then, we should show that $X_{t}$ is a solution to (2.5). For $t\in[0,T]$, we get
$\mathbb{E}|X_{t}-X^{n}_{t-1/n}|^{p}=$ $\mathbb{E}|X_{t}-X^{n}_{t}+X^{n}_{t}-X^{n}_{t-1/n}|^{p}$
$\leq$ $2^{p-1}\mathbb{E}|X_{t}-X^{n}_{t}|^{p}+2^{p-1}\mathbb{E}|X^{n}_{t}-X^{n}_{t-1/n}|^{p}.$
By Lemma 3.2 and (3.6), we get
$\mathbb{E}|X_{t}-X^{n}_{t-1/n}|^{p}\to 0,\ \ {\rm as}\ \ n\to\infty.$
Taking $n\to\infty$ in (3.1), we get
$X_{t}-D(X_{t-\tau})=\xi(0)+D(\xi(-\tau))+\int_{0}^{t}b(X_{s},X_{s-\tau},\mathcal{L}_{X_{s}}){\rm d}s+\int_{0}^{t}\sigma(\mathcal{L}_{X_{s}}){\rm d}B_{s}^{H},\ \ t\in[0,T].$
Step 2. Uniqueness. Let $X_{t}$ and $Y_{t}$ be two solutions of (2.5) with $X_{t}=Y_{t}=\xi,\ \ t\in[-\tau,0]$. By Lemma 2.7 again, we get
$|X_{t}-Y_{t}|^{p}\leq\lambda|X_{t-\tau}-Y_{t-\tau}|^{p}+\frac{1}{(1-\lambda)^{p-1}}|X_{t}-D(X_{t-\tau})+Y_{t}-D(Y_{t-\tau})|^{p},$
which implies,
$\sup_{0\leq r\leq t}|X_{r}-Y_{r}|^{p}$
$\leq$ $\sup_{-\tau\leq r\leq 0}|X_{r}-Y_{r}|^{p}+\lambda\sup_{0\leq r\leq t}|X_{r}-Y_{r}|^{p}+\frac{1}{(1-\lambda)^{p-1}}\sup_{0\leq r\leq t}|X_{r}-D(X_{r-\tau})+Y_{r}-D(Y_{r-\tau})|^{p}.$
Since $X_{t}=Y_{t}=\xi,\ \ t\in[-\tau,0]$ and $\lambda\in(0,1)$,
$\sup_{0\leq r\leq t}|X_{r}-Y_{r}|^{p}\leq\frac{1}{(1-\lambda)^{p}}\sup_{0\leq r\leq t}|X_{r}-D(X_{r-\tau})+Y_{r}-D(Y_{r-\tau})|^{p}=\frac{1}{(1-\lambda)^{p}}\sup_{0\leq r\leq t}|\Pi(r)|^{p},$
where $\Pi(r)=X_{r}-D(X_{r-\tau})+Y_{r}-D(Y_{r-\tau})$. For $t\in[0,T]$, by the Itô formula and the elementary inequality, we get
$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}|\Pi(s)|^{p}\bigg{)}\leq$ $p\mathbb{E}\int_{0}^{t}|\Pi(r)|^{p-2}\langle\Pi(r),b(X_{r},X_{r-\tau},\mathcal{L}_{X_{r}})-b(Y_{r},Y_{r-\tau},\mathcal{L}_{Y_{r}})\rangle{\rm d}r$
$+p(p-1)H\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}\int_{0}^{s}|\Pi(r)|^{p-2}|\sigma(\mathcal{L}_{X_{r}})-\sigma(\mathcal{L}_{Y_{r}})|^{2}{\rm d}r\bigg{)}$
$+p\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}\int_{0}^{s}|\Pi(r)|^{p-1}|\sigma(\mathcal{L}_{X_{r}})-\sigma(\mathcal{L}_{Y_{r}})|{\rm d}B_{r}^{H}\bigg{)}.$
$=$ $J_{1}+J_{2}+J_{3}.$
By Assumption 2.4, the Young’s inequality, we get
$J_{1}=$ $C\mathbb{E}\int_{0}^{t}|X_{r}-D(X_{r-\tau})+Y_{r}-D(Y_{r-\tau})|^{p}{\rm d}r$
$+C\mathbb{E}\int_{0}^{t}\bigg{[}|X_{r}-Y_{r}|^{2}+|X_{r-\tau}-Y_{r-\tau}|^{2}+(\mathbb{W}_{q}(\mathcal{L}_{X_{r}},\mathcal{L}_{Y_{r}}))^{2}\bigg{]}^{p/2}{\rm d}r$
$\leq$ $C\mathbb{E}\int_{0}^{t}\bigg{(}\sup_{0\leq u\leq r}|X_{u}-Y_{u}|^{p}\bigg{)}{\rm d}r+C\mathbb{E}\int_{0}^{t}\bigg{(}\sup_{0\leq u\leq r}|X_{u-\tau}-Y_{u-\tau}|^{p}\bigg{)}{\rm d}r.$
By Assumption 2.5 and the Young’s inequality, we get
$J_{2}\leq$ $C\mathbb{E}\int_{0}^{t}|X_{r}-D(X_{r-\tau})+Y_{r}-D(Y_{r-\tau})|^{p}{\rm d}r+C\mathbb{E}\int_{0}^{t}\bigg{[}(\mathbb{W}_{q}(\mathcal{L}_{X_{r}},\mathcal{L}_{Y_{r}}))^{2}\bigg{]}^{p/2}{\rm d}r$
$\leq$ $C\mathbb{E}\int_{0}^{t}\bigg{(}\sup_{0\leq u\leq r}|X_{u}-Y_{u}|^{p}\bigg{)}{\rm d}r+C\mathbb{E}\int_{0}^{t}\bigg{(}\sup_{0\leq u\leq r}|X_{u-\tau}-Y_{u-\tau}|^{p}\bigg{)}{\rm d}r.$
In the way in (3.3), by Assumption 2.5, the Young’s inequality, the Hölder inequality, we get
$J_{3}\leq$ $C\mathbb{E}\int_{0}^{t}\bigg{(}\sup_{0\leq u\leq r}|X_{u}-Y_{u}|^{p}\bigg{)}{\rm d}r+C\mathbb{E}\int_{0}^{t}\bigg{(}\sup_{0\leq u\leq r}|X_{u-\tau}-Y_{u-\tau}|^{p}\bigg{)}{\rm d}r.$
Combining $J_{1}$-$J_{3}$ and applying the Gronwall inequality, we get
$\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|X_{s}-Y_{s}|^{p}\bigg{)}\leq C\mathbb{E}\int_{0}^{t}\bigg{(}\sup_{0\leq u\leq r}|X_{u-\tau}-Y_{u-\tau}|^{p}\bigg{)}{\rm d}r.$
By induction, we derive that $X_{t}=Y_{t},t\in[0,T]$, $\mathbb{P}$-a.s. The proof is complete. ∎
4 Propagation of Chaos
Since the coefficients of (2.5) depend on the distributions of state variables, we use stochastic interacting particle systems to approximate it. Firstly, considering a non-interacting particle system is given by
(4.1) ${\rm d}(X_{t}^{i}-D(X_{t-\tau}^{i}))=b(X_{t}^{i},X_{t-\tau}^{i},\mathcal{L}_{X_{t}^{i}}){\rm d}t+\sigma(\mathcal{L}_{X_{t}^{i}}){\rm d}B_{t}^{H,i},\ \ X_{0}^{i}=Y_{0}^{i}=\xi,\ \ i=1,2,\cdots,N,$
here, $|\xi_{t}^{i}-\xi_{s}^{i}|^{p}\leq|t-s|^{p}$. To deal with $\mathcal{L}_{X_{t}^{i}}$, we exploit the following interacting particle system of the form
(4.2) ${\rm d}(X_{t}^{i,N}-D(X_{t-\tau}^{i,N}))=b(X_{t}^{i,N},X_{t-\tau}^{i,N},\mu_{t}^{X,N}){\rm d}t+\sigma(\mu_{t}^{X,N}){\rm d}B_{t}^{H,i},\ \ i=1,2,\cdots,N,$
where the empirical measures is defined by
$\mu^{X,N}_{t}(\cdot)=\frac{1}{N}\sum_{j=1}^{N}\delta_{X_{t}^{j,N}}(\cdot),$
here, $\delta_{x}$ denotes the Dirac measure at point $x$. In order to prove the desired result, we give the following Lemma.
Lemma 4.1.
[17, Lemma 3.2] For two empirical measures $\mu_{t}^{N}=\frac{1}{N}\sum\limits_{j=1}\limits^{N}\delta_{Z^{1,j,N}_{t}}$ and $\nu_{t}^{N}=\frac{1}{N}\sum\limits_{j=1}\limits^{N}\delta_{Z_{t}^{2,j,N}}$ , then
$\mathbb{E}\left(\mathbb{W}_{q}(\mu_{t}^{N},\nu_{t}^{N})\right)\leq\mathbb{E}\bigg{(}\frac{1}{N}\sum_{j=1}^{N}\bigg{\lvert}Z_{t}^{1,j,N}-Z_{t}^{2,j,N}\bigg{\rvert}^{q}\bigg{)}^{1/q}.$
Next, we consider the theory of the propagation of chaos.
Lemma 4.2.
Let Assumptions 2.2-2.5 hold. For $\overline{p}>1/H$, we have
$\mathbb{E}\bigg{(}\sup_{t\in[-\tau,T]}|X_{t}^{i}|^{\overline{p}}\bigg{)}+\mathbb{E}\bigg{(}\sup_{t\in[-\tau,T]}|X_{t}^{i,N}|^{\overline{p}}\bigg{)}\leq C,\ \ i=1,2,\cdots,N,$
where $C$ is a positive constant depending on ${\overline{p}},T,L,H,\xi$.
Proof.
We will proceed in the same way in Lemma 3.1. By Assumption 2.4 and Lemma 2.7, we get
$|X_{t}^{i,N}|^{\overline{p}}=|D(X_{t-\tau}^{i,N})+X_{t}^{n}-D(X_{t-\tau}^{i,N})|^{\overline{p}}\leq\lambda|X_{t-\tau}^{i,N}|^{\overline{p}}+\frac{1}{(1-\lambda)^{{\overline{p}}-1}}|X_{t}^{i,N}-D(X_{t-\tau}^{i,N})|^{\overline{p}},$
for any $t\in[0,T]$. We immediately get
$\sup_{0\leq r\leq t}|X_{r}^{i,N}|^{\overline{p}}\leq\frac{\lambda}{1-\lambda}\parallel\xi\parallel^{\overline{p}}+\frac{1}{(1-\lambda)^{{\overline{p}}}}\sup_{0\leq r\leq t}|X_{r}^{i,N}-D(X_{r-\tau}^{i,N})|^{\overline{p}}$
By the Itô formula, we get
$|X_{t}^{i,N}-D(X_{t-\tau}^{i,N})|^{\overline{p}}$
$\leq$ $|\xi(0)-D(\xi(-\tau))|^{\overline{p}}$
$+p\int_{0}^{t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{{\overline{p}}-2}\big{\langle}X_{s}^{i,N}-D(X_{s-\tau}),b\big{(}X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N}\big{)}\big{\rangle}{\rm d}s$
$+{\overline{p}}\int_{0}^{t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{p-1}\sigma\big{(}\mu_{s}^{X,N}\big{)}{\rm d}B_{s}^{H}$
$+{\overline{p}}({\overline{p}}-1)\int_{0}^{t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{{\overline{p}}-2}\sigma\big{(}\mu_{s}^{X,N}\big{)}D_{s}^{\phi}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|{\rm d}s$
=:absent:=:= : $|\xi(0)-D(\xi(-\tau))|^{\overline{p}}+I_{1}(t)+I_{2}(t)+I_{3}(t).$
Then,
$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}|X_{t}^{i,N}-D(X_{t-\tau}^{i,N})|^{\overline{p}}\bigg{)}\leq\mathbb{E}|\xi(0)-D(\xi(-\tau))|^{\overline{p}}+\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}I_{1}+I_{2}+I_{3}\bigg{)}$
By Assumptions 2.3, 2.4, and the Young’s inequality, we get
$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}I_{1}(s)\bigg{)}$
$=$ $\mathbb{E}\int_{0}^{t}{\overline{p}}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{{\overline{p}}-2}\big{\langle}X_{s}^{i,N}-D(X_{s-\tau}^{i,N}),b\big{(}X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N}\big{)}\big{\rangle}{\rm d}s$
$\leq$ $C{\overline{p}}\mathbb{E}\int_{0}^{t}[|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{p-2}\times(1+|X_{s}^{i,N}|^{2}+|X_{s-\tau}^{i,N}|^{2}+(\mathbb{W}_{q}(\mu_{s}^{X,N},\delta_{0})))^{2}]{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}[|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{\overline{p}}+(1+|X_{s}^{i,N}|^{2}+|X_{s-\tau}^{i,N}|^{2}+(\mathbb{W}_{q}(\mu_{s}^{X,N},\delta_{0}))^{2})^{{\overline{p}}/2}]{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t\wedge\nu_{R}}[(1+|X_{s}^{i,N}|^{\overline{p}}+|X_{s-\tau}^{i,N}|^{\overline{p}}+\big{(}\mathbb{W}_{q}(\mu_{s}^{X,N},\delta_{0})\big{)}^{{\overline{p}}}]{\rm d}s.$
By the Hölder inequality, Young’s inequality and Assumption 2.5 and in the way in (3.3), we get
$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}I_{2}(s)\bigg{)}\leq$ $C\mathbb{E}\int_{0}^{t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{{\overline{p}}-1}|\sigma(\mu_{s}^{X,N})|{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{\overline{p}}{\rm d}s+C\mathbb{E}\int_{0}^{t}((\mathbb{W}_{q}(\mu_{s}^{X,N},\delta_{0}))^{\overline{q}}+1){\rm d}s.$
By Assumption 2.5, Remark 2.1, the Young’s inequality, we get
$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}I_{3}(s)\bigg{)}$
$\leq$ $C\mathbb{E}\int_{0}^{t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{\overline{p}}{\rm d}s+C\mathbb{E}\int_{0}^{t}\bigg{[}|\sigma\big{(}\mu_{s}^{X,N}\big{)}|^{2}\int_{0}^{s}H(2H-1)|r-s|^{2H-2}{\rm d}r\bigg{]}^{{\overline{p}}/2}{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}|X_{s}^{i,N}-D(X_{s-\tau}^{i,N})|^{\overline{p}}{\rm d}s+2CHT^{2H-1}\mathbb{E}\int_{0}^{t}\lvert\sigma\big{(}\mu_{s}^{X,N}\big{)}\rvert^{\overline{p}}{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}[(|X_{s}^{n}|^{\overline{p}}+|X_{s-\tau}^{n}|^{\overline{p}}+\big{(}\mathbb{W}_{q}(\mu_{s}^{X,N},\delta_{0})\big{)}^{\overline{p}}]{\rm d}s$
$=$ $C\mathbb{E}\int_{0}^{t}[(|X_{s}^{n}|^{\overline{p}}+|X_{s-\tau}^{n}|^{\overline{p}}+C\mathbb{E}\int_{0}^{t}\bigg{(}\frac{1}{N}\sum_{j=1}^{N}|X_{r}^{j,N}|^{q}\bigg{)}^{{\overline{p}}/q}{\rm d}r,$
where for the last part, we uses the fact that
$\mathbb{W}_{\theta}(\mu_{s}^{X,N},\delta_{0})=\mathcal{W}_{q}(\mu_{s}^{X,N})=\bigg{(}\frac{1}{N}\sum_{j=1}^{N}|X_{s}^{j,N}|^{q}\bigg{)}^{1/q}.$
Since all $j$ are identically distributed, by the Minkowski inequality and Lemma 4.1, we get
$\mathbb{E}\bigg{(}\frac{1}{N}\sum_{j=1}^{N}|X_{s}^{j,N}|^{q}\bigg{)}^{{\overline{p}}/q}\leq\bigg{(}\frac{1}{N}\sum_{j=1}^{N}\parallel|X_{s}^{j,N}|^{q}\parallel_{L^{1/q}}\bigg{)}^{\overline{p}/q}=\bigg{(}\frac{1}{N}\sum_{j=1}^{N}(\mathbb{E}|X_{s}^{j,N}|)^{q}\bigg{)}^{\overline{p}/q}=\mathbb{E}|X_{s}^{i,N}|^{\overline{p}}.$
To sum up, we obtain
$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}|X_{t}^{i,N}-D(X_{t-\tau}^{i,N})|^{\overline{p}}\bigg{)}\leq$ $C+C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{0\leq u\leq s}|X_{u}^{i,N}|^{\overline{p}}\bigg{)}{\rm d}s+C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{0\leq u\leq s}|X_{u-\tau}^{i,N}|^{\overline{p}}\bigg{)}{\rm d}s,$
Thus, we get
$\mathbb{E}\bigg{(}\sup_{0\leq r\leq t}|X_{r}^{i,N}|^{\overline{p}}\bigg{)}\leq$ $C+C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{0\leq u\leq s}|X_{u}^{i,N}|^{\overline{p}}\bigg{)}{\rm d}s+C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{0\leq u\leq s}|X_{u-\tau}^{i,N}|^{\overline{p}}\bigg{)}{\rm d}s.$
For $t\in[0,\tau]$, we immediately get
$\mathbb{E}\bigg{(}\sup_{0\leq r\leq t}|X_{r}^{i,N}|^{\overline{p}}\bigg{)}\leq C+C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{0\leq u\leq s\wedge\nu_{R}}|X_{u}^{i,N}|^{\overline{p}}\bigg{)}{\rm d}s,$
since $\xi\in L_{\mathcal{F}_{0}}^{\overline{p}}([-\tau,0];\mathbb{R}^{d})$. We get the desired result by Gronwall inequality. By induction, we can derive the assertion. ∎
Lemma 4.3.
Let Assumption 2.4 hold. For $\overline{p}>1/H$ and $\overline{p}>2p$, we have
$\mathbb{E}\bigg{(}\sup_{t\in[0,T]}|X_{t}^{i}-X_{t}^{i,N}|^{p}\bigg{)}\leq C\times\left\{\begin{array}[]{ll}N^{-\frac{1}{2}},\ \ p>\frac{d}{2},\\ N^{-\frac{1}{2}}\log(1+N),\ p=\frac{d}{2},\\ N^{-\frac{p}{d}},\ p\in[2,\frac{d}{2}),\end{array}\right.$
where $C$ is a positive constant depending on $p,T,H,L,q$ but independent of $N$.
Proof.
For $i=1,2,\cdots,N$, by (4.2) and (4.1), set
$\Theta_{t}^{i}=X_{t}^{i}-D(X_{t-\tau}^{i})-X_{t}^{i,N}+D(X_{t-\tau}^{i,N}).$
By Itô formula, we get
$|\Theta_{t}^{i}|^{p}-|\Theta_{0}^{i}|^{p}$
$\leq$ $p\int_{0}^{t}|\Theta_{s}^{i}|^{p-1}\bigg{[}b(X_{s}^{i},X_{s-\tau}^{i},\mathcal{L}_{X_{s}^{i}})-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})\bigg{]}{\rm d}s$
$+p\int_{0}^{t}|\Theta_{t}^{i}|^{p-1}\bigg{[}\sigma(\mathcal{L}_{X_{s}^{i}})-\sigma(\mu_{s}^{X,N})\bigg{]}{\rm d}B_{s}^{H,i}$
$+p(p-1)\int_{0}^{t}|\Theta_{s}^{i}|^{p-2}\bigg{(}\sigma(\mathcal{L}_{X_{s}^{i}})-\sigma(\mu_{s}^{X,N})\bigg{)}\bigg{[}\int_{0}^{s}\phi(s,r)(\sigma(\mathcal{L}_{X_{r}^{i}})-\sigma(\mu_{r}^{X,N})){\rm d}r\bigg{]}{\rm d}s$
=:absent:=:= : $Q_{1}(t)+Q_{2}(t)+Q_{3}(t).$
For any $t\in[0,T]$, by Assumption 2.4, the Young’s inequality, we get
$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}Q_{1}(s)\bigg{)}$
$=$ $p\int_{0}^{t}|\Theta_{s}^{i}|^{p-1}\bigg{[}b(X_{s}^{i},X_{s-\tau}^{i},\mathcal{L}_{X_{s}^{i}})-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})\bigg{]}{\rm d}s$
$\leq$ $p\mathbb{E}\int_{0}^{t}|\Theta_{s}^{i}|^{p-2}\langle\Theta_{s}^{i},b(X_{s}^{i},X_{s-\tau}^{i},\mathcal{L}_{X_{s}^{i}})-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})\rangle{\rm d}s$
$\leq$ $C\int_{0}^{t}|\Theta_{s}^{i}|^{p}{\rm d}s+C\int_{0}^{t}\mathbb{E}[|X_{s}^{i}-X_{s}^{i,N}|^{p}+|X_{s-\tau}^{i}-X_{s-\tau}^{i,N}|^{p}+(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N}))^{p}]{\rm d}s.$
In the same way in (3.3), by Assumption 2.5, the Young’s inequality and the Hölder inequality, we get
$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}Q_{2}(s)\bigg{)}\leq$ $C\mathbb{E}\int_{0}^{t}|\Theta_{s}^{i}|^{\overline{p}-1}|\sigma(\mathcal{L}_{X_{s}^{i}})-\sigma(\mu_{s}^{X,N})|{\rm d}s$
$\leq$ $C\int_{0}^{t}|\Theta_{s}^{i}|^{p}{\rm d}s+C\int_{0}^{t}\mathbb{E}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N}))^{p}{\rm d}s.$
By Assumption 2.5, Remark 2.1, the Young’s inequality, we get
$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}Q_{3}(s)\bigg{)}\leq$ $C\mathbb{E}\int_{0}^{t}|\Theta_{s}^{i}|^{p}{\rm d}s+2CHT^{2H-1}\mathbb{E}\int_{0}^{t}\lvert\sigma\big{(}\mathcal{L}_{X_{u}^{i}}\big{)}-\sigma\big{(}\mu_{u}^{X,N}\big{)}\rvert^{p}{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}[(|\Theta_{s}^{i}|^{p}+\big{(}\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N})\big{)}^{p}]{\rm d}s.$
Noting that
$\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N})\leq\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X})+\mathbb{W}_{q}(\mu_{s}^{X},\mu_{s}^{X,N}),$
For $q<p$, by Lemma 4.1, we get
$\mathbb{E}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N}))^{p}=$ $\mathbb{E}[(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N}))^{q}]^{p/{q}}$
$\leq$ $\mathbb{E}\bigg{(}2^{q-1}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X}))^{q}+2^{q-1}(\mathbb{W}_{q}(\mu_{s}^{X},\mu_{s}^{X,N}))^{q}\bigg{)}^{p/q}$
$\leq$ $C\mathbb{E}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X}))^{p}+C\mathbb{E}(\mathbb{W}_{q}(\mu_{s}^{X},\mu_{s}^{X,N}))^{p}$
$\leq$ $C\mathbb{E}\bigg{(}\frac{1}{N}\sum_{j=1}^{N}|X_{s}^{j}-X_{s}^{j,N}|^{q}\bigg{)}^{p/q}+C\mathbb{E}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X}))^{p}$
$\leq$ $C\mathbb{E}|X_{s}^{i}-X_{s}^{i,N}|^{p}+C\mathbb{E}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X}))^{p}.$
Applying the Gronwall inequality, we get
$\mathbb{E}\bigg{(}\sup_{s\in[0,t]}|\Theta_{s}^{i}|^{p}\bigg{)}$
$\leq$ $C\int_{0}^{t}\mathbb{E}|X_{s}^{i}-X_{s}^{i,N}|^{p}{\rm d}s+C\int_{0}^{t}\mathbb{E}|X_{s-\tau}^{i}-X_{s-\tau}^{i,N}|^{p}{\rm d}s+C\int_{0}^{t}\mathbb{E}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N}))^{p}{\rm d}s$
$\leq$ $\mathbb{E}\int_{0}^{t}\bigg{(}\sup_{s\in[0,t]}|X_{s}^{i}-X_{s}^{i,N}|^{p}\bigg{)}{\rm d}s+C\int_{0}^{t}\mathbb{E}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X,N}))^{p}{\rm d}s$
$\leq$ $\mathbb{E}\int_{0}^{t}\bigg{(}\sup_{s\in[0,t]}|X_{s}^{i}-X_{s}^{i,N}|^{p}\bigg{)}{\rm d}s+C\int_{0}^{t}\mathbb{E}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X}))^{p}{\rm d}s.$
According to (2.7), we obtain
$|X_{s}^{i}-X_{s}^{i,N}|^{p}\leq\lambda|X_{t-\tau}^{i}-X_{t-\tau}^{i,N}|^{p}+\frac{1}{(1-\lambda)^{p-1}}|\Theta_{s}^{i}|^{p}.$
Noting that $10<\lambda<$ 1, by the Gronwall inequality, we get
$\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}|X_{s}^{i}-X_{s}^{i,N}|^{p}\bigg{)}\leq$ $C\mathbb{E}\bigg{(}\sup_{0\leq s\leq t}|\Theta_{s}^{i}|^{p}\bigg{)}\leq C\int_{0}^{t}\mathbb{E}(\mathbb{W}_{q}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X}))^{p}{\rm d}s.$
Since $(\mathbb{W}_{p}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X}))^{p}$ is controlled by the Wasserstein distance estimate([13], Theorem 1), for $\overline{p}>2p$, we get
$\mathbb{E}(\mathbb{W}_{p}(\mathcal{L}_{X_{s}^{i}},\mu_{s}^{X}))^{p}\leq$ $C(\mathbb{E}|X_{s}^{i}|^{\overline{p}})^{p/{\overline{p}}}\times\left\{\begin{array}[]{ll}N^{-\frac{1}{2}},\ \ p>\frac{d}{2},\\ N^{-\frac{1}{2}}\log(1+N),\ p=\frac{d}{2},\\ N^{-\frac{p}{d}},\ p\in[2,\frac{d}{2}).\end{array}\right.$
Then, by Lemma 4.2, we complete the proof. ∎
5 The tamed theta scheme
In this section, we consider the tamed theta scheme to approximate solution of (4.2). Since the drift coefficient exhibits super-linear growth, we define
(5.1) $b_{\Delta}(x,y,\mu)=\frac{b(x,y,\mu)}{1+\Delta^{\alpha}|b(x,y,\mu)|},$
for $x,y\in\mathbb{R}^{d}$, $\mu\in\mathcal{P}_{q}(\mathbb{R}^{d})$ and $\alpha\in(0,1/2]$. Assume that for any time $T>0$, there exist positive constants $M,m\in\mathbb{N}$ such that $\Delta=\frac{\tau}{m}=\frac{T}{M}$, where $\Delta\in(0,1)$ is the step-size. Consequently, for the time grid $t_{k}=t_{0}+k\Delta$, it holds that $t_{k+1}-\tau=t_{k+1-m}$. For $k=-m,\cdots,0,$ we set $Y_{t_{k}}^{i,N}=\xi^{i}(t_{k})$ and $t_{0}=0$. For $k=0,1,\cdots,M$, we now define the discrete EM scheme for (4.2) as follows,
(5.2) $Y_{t_{k+1}}^{i,N}-D(Y_{t_{k+1-m}}^{i,N})=$ $Y_{t_{k}}^{i,N}-D(Y_{t_{k-m}}^{i,N})+\theta b_{\Delta}(Y_{t_{k+1}}^{i,N},Y_{t_{k+1-m}}^{i,N},\mu_{t_{k+1}}^{Y,N})\Delta$
$+(1-\theta)b_{\Delta}(Y_{t_{k}}^{i,N},Y_{t_{k-m}}^{i,N},\mu_{t_{k}}^{Y,N})\Delta+\sigma(\mu_{t_{k}}^{Y,N})\Delta B_{t_{k}}^{H,i},$
where the empirical measures $\mu_{t_{k}}^{Y,N}(\cdot):=\frac{1}{N}\sum\limits_{j=1}^{N}\delta_{Y_{t_{k}}^{j,N}}(\cdot)$ and $\Delta B_{t_{k}}^{H,i}=B_{t_{k+1}}^{H,i}-B_{t_{k}}^{H,i}$. Let
(5.5) $\overline{Y}_{\Delta}^{i,N}(t)=\left\{\begin{array}[]{ll}\xi^{i}(t),\ \ -\tau\leq t\leq 0,\\ \sum\limits_{k=0}^{[T/\Delta]}Y_{t_{k}}^{i,N}{\rm I}_{[k\Delta,(k+1)\Delta)}(t),\ \ t\geq 0,\end{array}\right.$
and
$\overline{Y}_{\Delta+}^{i,N}(t)=\left\{\begin{array}[]{ll}\xi^{i}(t),\ \ -\tau\leq t\leq 0,\\ \sum\limits_{k=0}^{[T/\Delta]}Y_{t_{k+1}}^{i,N}{\rm I}_{[k\Delta,(k+1)\Delta)}(t),\ \ t\geq 0,\end{array}\right.$
and the continuous EM solution writes as
(5.6) $Y_{\Delta}^{i,N}(t)=$ $D(\overline{Y}_{\Delta}^{i,N}(t))+\xi^{i}(0)-D(\xi^{i}(-\tau))+\theta\int_{0}^{t}b_{\Delta}(\overline{Y}_{\Delta+}^{i,N}(s),\overline{Y}_{\Delta+}^{i,N}(s-\tau),\overline{\mu}^{Y_{+},N}(s)){\rm d}s$
$+(1-\theta)\int_{0}^{t}b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s)){\rm d}s+\int_{0}^{t}\sigma(\overline{\mu}^{Y,N}(s)){\rm d}B_{s}^{H,i},\ \ t>0,$
where $\overline{\mu}_{s}^{Y,N}:=\frac{1}{N}\sum\limits_{j=1}^{N}\delta_{\overline{Y}_{t}^{j,N}}(\cdot)$, and $Y_{\Delta}^{i,N}(t)=\xi^{i}(t)$ when $t_{k}\in[-\tau,0]$. However, this $Y_{\Delta}^{i,N}(t)$ is not $\mathcal{F}_{t}$-adapted, it does not meet the fundamental requirement in the Itô stochastic analysis. To avoid calculus, we use the discrete split-step theta scheme (see [28]) described as follows: For $k=-m,\cdots,-1$, set $z_{t_{k}}^{i,N}=\xi^{i}(k\Delta)$. For $k=0,1,\cdots,M-1$, we rewrite (5.2) as
(5.9) $\left\{\begin{array}[]{ll}Y_{t_{k}}^{i,N}=D(Y_{t_{k-m}}^{i,N})+z_{t_{k}}^{i,N}-D(z_{t_{k-m}})+\theta b_{\Delta}(Y_{t_{k}}^{i,N},Y_{t_{k-m}}^{i,N},\mu_{t_{k}}^{Y,N})\Delta,\\ z_{t_{k+1}}^{i,N}=D(z_{t_{k+1-m}}^{i,N})+z_{t_{k}}^{i,N}-D(z_{t_{k-m}}^{i,N})+b_{\Delta}(Y_{t_{k}}^{i,N},Y_{t_{k-m}}^{i,N},\mu_{t_{k}}^{Y,N})\Delta+\sigma(\mu_{t_{k}}^{Y,N})\Delta B_{t_{k}}^{H,i},\end{array}\right.$
This scheme can also be rewritten as
$z_{t_{k+1}}^{i,N}-D(z_{t_{k+1-m}}^{i,N})=z_{t_{0}}^{i,N}-D(z_{t_{-m}}^{i,N})+\sum_{j=0}^{k}b_{\Delta}(Y_{t_{j}}^{i,N},Y_{t_{j-m}}^{i,N},\mu_{t_{j}}^{Y,N})\Delta+\sum_{j=0}^{k}\sigma(\mu_{t_{j}}^{Y,N})\Delta B_{t_{j}}^{H,i}$
$=$ $\xi^{i}(0)-D(\xi^{i}(-\tau))-\theta b_{\Delta}(\xi^{i}(0),\xi^{i}(-\tau),\mu_{t_{0}}^{i,N})\Delta+\sum_{j=0}^{k}b_{\Delta}(Y_{t_{j}}^{i,N},Y_{t_{j-m}}^{i,N},\mu_{t_{j}}^{Y,N})\Delta$
$+\sum_{j=0}^{k}\sigma(\mu_{t_{j}}^{Y,N})\Delta B_{t_{j}}^{H,i}.$
Similarly, we define the corresponding continuous-time split-step tamed theta solution $Z_{\Delta}^{i,N}(t)$ as follows: For any $t\in[-\tau,0]$, $Z_{\Delta}^{i,N}(t)=\xi^{i}(t)$, $Z_{\Delta}^{i,N}(0)=\xi^{i}(0)-\theta b_{\Delta}(\xi^{i}(0)$, $\xi^{i}(-\tau),\mu_{0}^{Y,N})\Delta$. For any $t\in[0,T]$,
(5.10) ${\rm d}[Z_{\Delta}^{i,N}(t)-D(Z_{\Delta}^{i,N}(t-\tau))]=b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(t),\overline{Y}_{\Delta}^{i,N}(t-\tau),\overline{\mu}^{Y,N}(t)){\rm d}t+\sigma(\overline{\mu}^{Y,N}(t)){\rm d}B_{t}^{H,i},$
where $\overline{Y}_{\Delta}^{i,N}(t)$ is defined by (5.5). Combining (5.6), (5.9) and (5.10), we have
$Y_{\Delta}^{i,N}(t)-D(Y_{\Delta}^{i,N}(t-\tau))-\theta b_{\Delta}(Y_{\Delta}^{i,N}(t),Y_{\Delta}^{i,N}(t-\tau),\mu^{Y,N}(t))\Delta=Z_{\Delta}^{i,N}(t)-D(Z_{\Delta}^{i,N}(t-\tau)).$
Let $\widetilde{Y}_{\Delta}^{i,N}(t)=Y_{\Delta}^{i,N}(t)-D(Y_{\Delta}^{i,N}(t-\tau))-\theta b_{\Delta}(Y_{\Delta}^{i,N}(t),Y_{\Delta}^{i,N}(t-\tau),\mu^{Y,N}(t))\Delta$. We can rewrite (5.10) as
(5.11) $\widetilde{Y}_{\Delta}^{i,N}(t)=\widetilde{Y}_{\Delta}^{i,N}(0)+\int_{0}^{t}b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s)){\rm d}s+\int_{0}^{t}\sigma(\overline{\mu}^{Y,N}(s)){\rm d}B_{s}^{H,i}.$
It is clear that $\widetilde{Y}_{\Delta}^{i,N}$ coincide with $\overline{Y}_{\Delta}^{i,N}(t)-D(\overline{Y}_{\Delta}^{i,N}(t-\tau))-\theta b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(t),\overline{Y}_{\Delta}^{i,N}(t-\tau),\overline{\mu}^{Y,N}(t))$ at grid points $t=k\Delta$, $k=0,1,\cdots,M-1$, this also means that the continuous-time tamed theta solution $Y_{\Delta}^{i,N}(t)$ and the discrete-time tamed theta solution $\overline{Y}_{\Delta}^{i,N}(t)$ take the same values at grid points $t=k\Delta$, $k=0,1,\cdots,M-1$.
Lemma 5.1.
There exists a positive constant $K_{7}\geq 1$ such that
$|b_{\Delta}(x,y,\mu)|\leq\min\{K_{7}\Delta^{-\alpha}(1+|x|+|y|+\mathbb{W}_{q}(\mu,\delta_{0})),|b(x,y,\mu)|\}.$
Proof.
For $\forall x,y>0$, by (5.1), it is easy to see that $|b_{\Delta}(x,y,\mu)|\leq|b(x,y,\mu)|$. On the other hand, we immediately get
$|b_{\Delta}(x,y,\mu)|=\Delta^{-\alpha}\frac{|b(x,y,\mu)|}{\Delta^{-\alpha}+|b(x,y,\mu)|}\leq\Delta^{-\alpha}\leq K_{7}\Delta^{-\alpha}(1+|x|+|y|+\mathbb{W}_{q}(\mu,\delta_{0})).$
Lemma 5.2.
Let Assumption 2.4 hold. One observes
$\langle x-D(y)-\overline{x}+D(\overline{y}),b_{\Delta}(x,y,\mu)-b_{\Delta}(\overline{x},\overline{y},\nu)\rangle\leq K_{2}(|x-\overline{x}|^{2}+|y-\overline{y}|^{2}+(\mathbb{W}_{q}(\mu,\nu))^{2}),$
and
$|b(x,y,\mu)-b_{\Delta}(x,y,\mu)|^{p}\leq K_{8}\Delta^{\alpha p}(1+|x|^{2(l+1)p}+|y|^{2(l+1)p}+(\mathbb{W}_{q}(\mu,\delta_{0}))^{2p}),$
where $K_{8}$ is a positive constant.
Proof.
We divide it into two case. Case 1, for $b(x,y,\mu)\cdot b(\overline{x},\overline{y},\nu)<0$, we get
$\langle x-D(y)-\overline{x}+D(\overline{y}),b_{\Delta}(x,y,\mu)-b_{\Delta}(\overline{x},\overline{y},\nu)\rangle$
$=$ $\bigg{\langle}x-D(y)-\overline{x}+D(\overline{y}),\frac{b(x,y,\mu)}{1+\Delta^{\alpha}|b(x,y,\mu)|}-\frac{b(\overline{x},\overline{y},\nu)}{1+\Delta^{\alpha}|b(\overline{x},\overline{y},\nu)|}\bigg{\rangle}$
$\leq$ $K_{2}(|x-\overline{x}|^{2}+|y-\overline{y}|^{2}+(\mathbb{W}_{q}(\mu,\nu))^{2}).$
Case 2, for $b(x,y,\mu)\cdot b(\overline{x},\overline{y},\nu)>0$, we get
$\langle x-D(y)-\overline{x}+D(\overline{y}),b_{\Delta}(x,y,\mu)-b_{\Delta}(\overline{x},\overline{y},\nu)\rangle$
$=$ $\bigg{\langle}x-D(y)-\overline{x}+D(\overline{y}),\frac{b(x,y,\mu)-b(\overline{x},\overline{y},\nu)}{(1+\Delta^{\alpha}|b(x,y,\mu)|)(1+\Delta^{\alpha}|b(\overline{x},\overline{y},\nu))}\bigg{\rangle}$
$+\bigg{\langle}x-D(y)-\overline{x}+D(\overline{y}),\frac{\Delta^{\alpha}[b(x,y,\mu)\cdot b(\overline{x},\overline{y},\nu)-b(x,y,\mu)\cdot b(\overline{x},\overline{y},\nu)]}{(1+\Delta^{\alpha}|b(x,y,\mu)|)(1+\Delta^{\alpha}|b(\overline{x},\overline{y},\nu))}\bigg{\rangle}$
$\leq$ $K_{2}(|x-\overline{x}|^{2}+|y-\overline{y}|^{2}+(\mathbb{W}_{q}(\mu,\nu))^{2}).$
Furthermore,
$|b(x,y,\mu)-b_{\Delta}(x,y,\mu)|^{p}\leq\Delta^{\alpha p}|b(x,y,\mu)|^{2p},$
the second part is obvious. ∎
Remark 5.3.
Given Lemmas 5.1, 5.2 hold. We have
$\langle x-D(y),b_{\Delta}(x,y,\mu)\rangle\leq C(1+|x|^{2}+|y|^{2}+(\mathbb{W}_{q}(\mu,\delta_{0}))^{2}).$
5.1 Moment bounds
In order to get the desired results, we give the following lemma.
Lemma 5.4.
Let Assumption 2.2-2.5 and Lemmas 5.1, 5.2 hold. For $\overline{p}>1/H$, we have
$\sup_{0\leq t\leq T}\mathbb{E}|Y_{\Delta}^{i,N}(t)|^{\overline{p}}\leq C,$
where the positive constant $C$ is independent of $\Delta$.
Proof.
For $a>0$, let $\lfloor a\rfloor$ be the integer part of $a$. Applying the Itô formula to $[1+|\widetilde{Y}_{\Delta}^{i,N}(t)|^{2}]^{\overline{p}/2}$, we get
$\mathbb{E}[1+|\widetilde{Y}_{\Delta}^{i,N}(t)|^{2}]^{\overline{p}/2}$
$\leq$ $\mathbb{E}[1+|\widetilde{Y}_{\Delta}^{i,N}(0)|^{2}]^{\overline{p}/2}+p\mathbb{E}\int_{0}^{t}[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\rangle$
$+\overline{p}(\overline{p}-2)\mathbb{E}\int_{0}^{t}[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}\sigma(\overline{\mu}^{Y,N}(s))D_{s}^{\phi}[1+|\widetilde{Y}_{\Delta}(s)|^{2}]{\rm d}s$
$\leq$ $\mathbb{E}[1+|\widetilde{Y}_{\Delta}^{i,N}(0)|^{2}]^{\overline{p}/2}+\overline{p}(\overline{p}-2)\mathbb{E}\int_{0}^{t}[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}|\sigma(\overline{\mu}^{Y,N}(s))|^{2}\int_{0}^{s}\phi(s,r){\rm d}r{\rm d}s$
$+\overline{p}\mathbb{E}\int_{0}^{t}[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}\langle\overline{Y}_{\Delta}^{i,N}(s)-D(\overline{Y}_{\Delta}^{i,N}(s-\tau)),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\rangle$
$+\overline{p}\mathbb{E}\int_{0}^{t}[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(s)-\widehat{Y}_{\Delta}^{i,N}(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\rangle{\rm d}s$
=:absent:=:= : $\mathbb{E}[1+|\widetilde{Y}_{\Delta}^{i,N}(0)|^{2}]^{\overline{p}/2}+G_{1}(t)+G_{2}(t)+G_{3}(t).$
where $\widehat{Y}_{\Delta}^{i,N}(s)=\overline{Y}_{\Delta}^{i,N}(s)-D(\overline{Y}_{\Delta}^{i,N}(s-\tau))-\theta b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\Delta$. By the properties of $\phi$ and the Young’s inequality, Lemma 5.1, we get
$G_{1}(t)$
$\leq$ $CHT^{2H-1}\mathbb{E}\int_{0}^{t}[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}|\sigma(\overline{\mu}^{Y,N}(s))|^{2}{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}\big{(}1+[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}}{2}}+|\overline{Y}_{\Delta}^{i,N}(s)|^{\overline{p}}\big{)}{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}\bigg{(}1+|Y_{\Delta}^{i,N}(s)|^{\overline{p}}+|Y_{\Delta}^{i,N}(s-\tau)|^{\overline{p}}+|\theta b_{\Delta}(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))\Delta|^{\overline{p}}+|\overline{Y}_{\Delta}^{i,N}(s)|^{\overline{p}}\bigg{)}{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}\bigg{(}1+|Y_{\Delta}^{i,N}(s)|^{\overline{p}}+|Y_{\Delta}^{i,N}(s-\tau)|^{\overline{p}}+|\overline{Y}_{\Delta}^{i,N}(s)|^{\overline{p}}\bigg{)}{\rm d}s$
$+C\Delta^{(1-\alpha)\overline{p}}\mathbb{E}\int_{0}^{t}\bigg{(}1+|Y_{\Delta}^{i,N}(s)|^{\overline{p}}+|Y_{\Delta}^{i,N}(s-\tau)|^{\overline{p}}\bigg{)}$
$\leq$ $C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)|^{\overline{p}}{\rm d}s.$
By Remark 5.3 and the Young’s inequality, we get
$G_{2}(t)\leq$ $C\mathbb{E}\int_{0}^{t}[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}\bigg{(}1+|\overline{Y}_{\Delta}^{i,N}(s)|^{2}+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|^{2}+(\mathbb{W}_{q}(\mu^{\overline{Y},N}(s),\delta_{0}))^{2}\bigg{)}{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}\bigg{(}[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}}{2}}+1+|\overline{Y}_{\Delta}^{i,N}(s)|^{\overline{p}}+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|^{\overline{p}}\bigg{)}{\rm d}s$
$\leq$ $C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)|^{\overline{p}}{\rm d}s.$
We may compute that
$G_{3}(t)=$ $\overline{p}\mathbb{E}\int_{0}^{t}[1+|\widehat{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(s)-\widehat{Y}_{\Delta}^{i,N}(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\rangle{\rm d}s$
$+p\mathbb{E}\int_{0}^{t}\bigg{[}[1+|\widetilde{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}-[1+|\widehat{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}\bigg{]}$
×$\times\langle\widetilde{Y}_{\Delta}^{i,N}(s)-\widehat{Y}_{\Delta}^{i,N}(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\rangle{\rm d}s$
=:absent:=:= : $\overline{p}G_{31}(t)+\overline{p}G_{32}(t),$
where $\widetilde{Y}_{\Delta}^{i,N}(t)-\widehat{Y}_{\Delta}^{i,N}(s)=\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u)){\rm d}u+\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}\sigma(\overline{\mu}^{Y,N}(u)){\rm d}B_{u}^{H,i}$. By Lemma 5.1 and the Young’s inequality, we get
$G_{31}(t)$
$=$ $\mathbb{E}\int_{0}^{t}[1+|\widehat{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}\bigg{\langle}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(u)){\rm d}u,$
$b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\bigg{\rangle}{\rm d}s$
$+\mathbb{E}\int_{0}^{t}[1+|\widehat{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}\bigg{\langle}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}\sigma(\overline{\mu}^{Y,N}(u)){\rm d}B_{u}^{H,i},b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\bigg{\rangle}{\rm d}s$
$\leq$ $\mathbb{E}\int_{0}^{t}[1+|\widehat{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(u))|{\rm d}u$
×$\times|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))|{\rm d}s$
$\leq$ $\Delta\mathbb{E}\int_{0}^{t}[1+|\widehat{Y}_{\Delta}^{i,N}(s)|^{2}]^{\frac{\overline{p}-2}{2}}|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))|^{2}{\rm d}s$
$\leq$ $C\Delta^{1-2\alpha}\mathbb{E}\int_{0}^{t}\bigg{[}(1+|\overline{Y}_{\Delta}^{i,N}(u)|^{\overline{p}}+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|^{\overline{p}})+\Delta^{(1-\alpha)\overline{p}}(1+|\overline{Y}_{\Delta}^{i,N}(u)|^{\overline{p}}+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|^{\overline{p}})\bigg{]}{\rm d}s$
$\leq$ $C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)|^{\overline{p}}{\rm d}s.$
Applying Itô formula to $[1+|\widetilde{Y}_{\Delta}(s)|^{2}]^{\frac{\overline{p}-2}{2}}$, we get
$[1+|\widetilde{Y}_{\Delta}(s)|^{2}]^{\frac{\overline{p}-2}{2}}$
$\leq$ $[1+|\widetilde{Y}_{\Delta}^{i,N}(0)|^{2}]^{\frac{\overline{p}-2}{2}}$
$+(p-2)\int_{0}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u))\rangle{\rm d}u$
$+(\overline{p}-2)(\overline{p}-4)\int_{0}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\sigma(\overline{\mu}^{Y,N}(u))D_{r}^{\phi}[1+|\widetilde{Y}_{\Delta}(u)|^{2}]^{\frac{\overline{p}-2}{2}}{\rm d}u$
$+(\overline{p}-2)\int_{0}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),\sigma(\overline{\mu}^{Y,N}(u))\rangle{\rm d}B_{u}^{H,i}.$
Then,
$[1+|\widehat{Y}_{\Delta}(s)|^{2}]^{\frac{\overline{p}-2}{2}}$
$\leq$ $[1+|\widehat{Y}_{\Delta}^{i,N}(0)|^{2}]^{\frac{\overline{p}-2}{2}}$
$+(\overline{p}-2)\int_{0}^{\lfloor\frac{s}{\Delta}\rfloor\Delta}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u))\rangle{\rm d}u$
$+(\overline{p}-2)(\overline{p}-4)\int_{0}^{\lfloor\frac{s}{\Delta}\rfloor\Delta}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\sigma(\overline{\mu}^{Y,N}(u))D_{u}^{\phi}[1+|\widetilde{Y}_{\Delta}(u)|^{2}]^{\frac{\overline{p}-2}{2}}{\rm d}u$
$+(\overline{p}-2)\int_{0}^{\lfloor\frac{s}{\Delta}\rfloor\Delta}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),\sigma(\overline{\mu}^{Y,N}(u))\rangle{\rm d}B_{u}^{H,i}.$
One immediately gets
$G_{32}(t)\leq$ $(\overline{p}-2)\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u))\rangle{\rm d}u$
×$\times\langle\widetilde{Y}_{\Delta}^{i,N}(s)-\widehat{Y}_{\Delta}^{i,N}(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\rangle{\rm d}s$
$+(\overline{p}-2)(\overline{p}-4)\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\sigma(\overline{\mu}^{Y,N}(u))D_{u}^{\phi}[1+|\widetilde{Y}_{\Delta}(u)|^{2}]^{\frac{\overline{p}-2}{2}}{\rm d}u$
×$\times\langle\widetilde{Y}_{\Delta}^{i,N}(s)-\widehat{Y}_{\Delta}^{i,N}(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\rangle{\rm d}s$
$+(\overline{p}-2)\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),\sigma(\overline{\mu}^{Y,N}(u))\rangle{\rm d}B_{u}^{H,i}$
×$\times\langle\widetilde{Y}_{\Delta}^{i,N}(s)-\widehat{Y}_{\Delta}^{i,N}(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\rangle{\rm d}s$
=:absent:=:= : $(\overline{p}-2)G_{321}(t)+(\overline{p}-2)(\overline{p}-4)G_{322}(t)+(\overline{p}-2)G_{323}(t).$
By Lemma 5.1, the Young’s inequality, we get
$G_{321}(t)\leq$ $\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u))\rangle{\rm d}u$
×$\times\bigg{\langle}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u)){\rm d}u,b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\bigg{\rangle}{\rm d}s$
$+\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u))\rangle{\rm d}u$
×$\times\bigg{\langle}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}\sigma(\overline{\mu}^{Y,N}(u)){\rm d}B_{u}^{H,i},b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\bigg{\rangle}{\rm d}s$
$\leq$ $\Delta\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-3}{2}}|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))|^{3}{\rm d}u{\rm d}s$
$+C\mathbb{E}\int_{0}^{t}\bigg{[}\bigg{(}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-3}{2}}|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u),\overline{\mu}^{Y,N}(u))|{\rm d}u$
×$\times|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))|\bigg{)}^{\frac{\overline{p}}{\overline{p}-1}}+\bigg{\lvert}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}\sigma(\overline{\mu}^{Y,N}(u)){\rm d}B_{u}^{H,i}\bigg{\rvert}^{\overline{p}}\bigg{]}{\rm d}s$
$\leq$ $C\Delta^{2-3\alpha}\mathbb{E}\int_{0}^{t}(1+|Y_{\Delta}^{i,N}(s)|^{\overline{p}}+|Y_{\Delta}^{i,N}(s-\tau)|^{\overline{p}}+|\overline{Y}_{\Delta}^{i,N}(s)|^{p}+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|^{\overline{p}}){\rm d}s$
$+C\Delta^{2-3\alpha}\Delta^{(1-\alpha)\overline{p}}\mathbb{E}\int_{0}^{t}(1+|Y_{\Delta}^{i,N}(s)|^{\overline{p}}+|Y_{\Delta}^{i,N}(s-\tau)|^{p}+|\overline{Y}_{\Delta}^{i,N}(s)|^{\overline{p}}$
$+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|^{\overline{p}}){\rm d}s$
$+C\mathbb{E}\int_{0}^{t}\bigg{[}\bigg{(}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-3}{2}}|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))|^{2}{\rm d}u|\bigg{)}^{\frac{\overline{p}}{\overline{p}-1}}$
$+\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}\lvert\sigma(\overline{\mu}^{Y,N}(u))\rvert^{\overline{p}}{\rm d}u{\rm d}s$
$\leq$ $C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(s)|^{\overline{p}}{\rm d}s.$
Similarly, we get
$G_{322}(t)\leq C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(s)|^{\overline{p}}{\rm d}s.$
By Assumption 2.4, Lemma 5.1 and [32, Lemma 3.1.3], noting that $\mathbb{E}\int_{0}^{t}F_{s}{\rm d}B_{s}^{H,i}=0$, we get
$G_{323}(t)=$ $\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),\sigma(\overline{\mu}^{Y,N}(u))\rangle{\rm d}B_{u}^{H,i}$
×$\times\bigg{\langle}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(u),\overline{Y}_{\Delta}^{i,N}(u-\tau),\overline{\mu}^{Y,N}(u)){\rm d}u,b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\bigg{\rangle}{\rm d}s$
$+\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),\sigma(\overline{\mu}^{Y,N}(u))\rangle{\rm d}B_{u}^{H,i}$
×$\times\bigg{\langle}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}\sigma(\overline{\mu}^{Y,N}(u)){\rm d}B_{u}^{H,i},b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\bigg{\rangle}{\rm d}s$
$=$ $\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-4}{2}}\langle\widetilde{Y}_{\Delta}^{i,N}(u),\sigma(\overline{\mu}^{Y,N}(u))\rangle{\rm d}B_{u}^{H,i}$
×$\times\bigg{\langle}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}\sigma(\overline{\mu}^{Y,N}(u)){\rm d}B_{u}^{H,i},b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))\bigg{\rangle}{\rm d}s$
$\leq$ $2HT^{2H-2}\mathbb{E}\int_{0}^{t}\int_{\lfloor\frac{s}{\Delta}\rfloor\Delta}^{s}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\frac{\overline{p}-3}{2}}|\sigma(\overline{\mu}^{Y,N}(u))|^{2}{\rm d}u$
×$\times|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s),\overline{\mu}^{Y,N}(s))|{\rm d}s$
$\leq$ $C\Delta^{1-\alpha}\mathbb{E}\int_{0}^{t}(|Y_{\Delta}^{i,N}(s)|^{\overline{p}}+|Y_{\Delta}^{i,N}(s-\tau)|^{\overline{p}}+|\overline{Y}_{\Delta}^{i,N}(s))|^{\overline{p}}){\rm d}s$
$+C\Delta^{1-\alpha}\Delta^{(1-\alpha)\overline{p}}\mathbb{E}\int_{0}^{t}|\overline{Y}_{\Delta}^{i,N}(s))|^{\overline{p}}{\rm d}s$
$\leq$ $C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(s)|^{\overline{p}}{\rm d}s.$
To sum up, we derive
$G_{3}(t)\leq C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(s)|^{\overline{p}}{\rm d}s.$
Combining $G_{1}-G_{3}$, we get
(5.12) $\sup_{0\leq u\leq t}\mathbb{E}|\widetilde{Y}_{\Delta}^{i,N}(u)|^{\overline{p}}\leq\sup_{0\leq u\leq t}\mathbb{E}[1+|\widetilde{Y}_{\Delta}^{i,N}(u)|^{2}]^{\overline{p}/2}\leq C+C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)|^{\overline{p}}{\rm d}s.$
Noting that $|x-y|^{\overline{p}}\geq 2^{1-\overline{p}}|x|^{\overline{p}}-|y|^{\overline{p}}$, by Assumption 2.3, Lemma 5.1, we get
$|\widetilde{Y}_{\Delta}^{i,N}(u)|^{\overline{p}}\geq$ $2^{1-\overline{p}}|Y_{\Delta}^{i,N}(t)-D(Y_{\Delta}^{i,N}(t-\tau))|^{\overline{p}}-|\theta b_{\Delta}(Y_{\Delta}^{i,N}(t),Y_{\Delta}^{i,N}(t-\tau),\mu^{Y,N}(t))\Delta|^{\overline{p}}$
$\geq$ $2^{1-\overline{p}}[2^{1-\overline{p}}|Y_{\Delta}^{i,N}(t)|^{\overline{p}}-|D(Y_{\Delta}^{i,N}(t-\tau))|^{\overline{p}}]-|\theta b_{\Delta}(Y_{\Delta}^{i,N}(t),Y_{\Delta}^{i,N}(t-\tau),\mu^{Y,N}(t))\Delta|^{\overline{p}}$
$\geq$ $(2^{2-2\overline{p}}-4^{\overline{p}-1}\theta^{\overline{p}}K_{7}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}})|Y_{\Delta}^{i,N}(t)|^{\overline{p}}-(2^{1-\overline{p}}\lambda^{\overline{p}}+4^{\overline{p}-1}\theta^{p}K_{7}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}})|Y_{\Delta}^{i,N}(t-\tau)|^{\overline{p}}$
$-4^{\overline{p}-1}\theta^{\overline{p}}K_{7}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}}(\mathbb{W}_{q}(\mu^{Y,N}(t)),\delta_{0})^{\overline{p}}-4^{\overline{p}-1}\theta^{\overline{p}}K_{7}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}}.$
Thus, we obtain
$\sup_{0\leq u\leq t}\mathbb{E}|Y_{\Delta}^{i,N}(t)|^{\overline{p}}\leq(2^{2-2\overline{p}}-3\times 4^{\overline{p}-1}\theta^{\overline{p}}K_{7}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}}-2^{1-\overline{p}}\lambda^{\overline{p}})^{-1}$
×$\times\bigg{[}\sup_{0\leq u\leq t}\mathbb{E}|\widetilde{Y}_{\Delta}^{i,N}(u)|^{\overline{p}}+4^{\overline{p}-1}\theta^{\overline{p}}K_{7}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}}+(2^{1-\overline{p}}\lambda^{\overline{p}}+4^{\overline{p}-1}\theta^{\overline{p}}K_{7}^{\overline{p}}\Delta^{(1-\alpha)\overline{p}})\mathbb{E}\parallel\xi\parallel^{\overline{p}}\bigg{]}.$
Then, we derive from (5.12),
$\sup_{0\leq u\leq t}\mathbb{E}|Y_{\Delta}^{i,N}(t)|^{\overline{p}}\leq C+C\int_{0}^{t}\sup_{0\leq u\leq t}\mathbb{E}|Y_{\Delta}^{i,N}(t)|^{\overline{p}}.$
Applying the Gronwall inequality, we complete the proof. ∎
5.2 Strong convergence rate
Lemma 5.5.
Let Assumptions 2.2-2.5 hold. For $p>1/H$, then
$\mathbb{E}\bigg{(}\sup_{0\leq k\leq M-1}\sup_{t_{k}\leq t\leq t_{k+1}}|Y_{\Delta}^{i,N}(t)-Y_{\Delta}^{i,N}(t_{k})|^{p}\bigg{)}\leq C\Delta^{[(1-\alpha)\wedge H]p},$
where $C$ is a positive constant independent of $\Delta$.
Proof.
For $t\in[t_{k},t_{k+1}]$, by the elementary inequality, we may compute
$\mathbb{E}\bigg{(}\sup_{t_{k}\leq t\leq t_{k+1}}|\widetilde{Y}_{\Delta}^{i,N}(t)-\widetilde{Y}_{\Delta}^{i,N}(t_{k})|^{p}\bigg{)}$
$=$ $\mathbb{E}\bigg{(}\sup_{t_{k}\leq t\leq t_{k+1}}\bigg{\lvert}\int_{t_{k}}^{t}b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s)){\rm d}s+\int_{t_{k}}^{t}\sigma(\overline{\mu}^{Y,N}(s)){\rm d}B_{s}^{H,i}\bigg{\rvert}^{p}\bigg{)}$
$\leq$ $2^{p-1}\mathbb{E}\bigg{(}\sup_{t_{k}\leq t\leq t_{k+1}}\bigg{\lvert}\int_{t_{k}}^{t}b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s)){\rm d}s\bigg{\rvert}^{p}\bigg{)}$
$+2^{p-1}\mathbb{E}\bigg{(}\sup_{t_{k}\leq t\leq t_{k+1}}\bigg{\lvert}\int_{t_{k}}^{t}\sigma(\overline{\mu}^{Y,N}(s)){\rm d}B_{s}^{H,i}\bigg{\rvert}^{p}\bigg{)}$
=:absent:=:= : $J_{1}+J_{2}.$
By the Hölder inequality and Lemmas 5.1, 5.4, we get
$J_{1}\leq$ $2^{p-1}\Delta^{p-1}\mathbb{E}\int_{t_{k}}^{t_{k+1}}|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))|^{p}{\rm d}s$
$\leq$ $2^{p-1}\Delta^{p-1}\mathbb{E}\int_{t_{k}}^{t_{k+1}}K_{6}\Delta^{-\alpha p}(1+|\overline{Y}_{\Delta}^{i,N}(s)|+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|+\mathbb{W}_{q}(\overline{\mu}^{Y,N}(s),\delta_{0}))^{p}{\rm d}s$
$\leq$ $C\Delta^{(1-\alpha)p}+C\Delta^{(1-\alpha)p}\mathbb{E}\int_{t_{k}}^{t_{k+1}}(|\overline{Y}_{\Delta}^{i,N}(s)|^{p}+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|^{p}){\rm d}s$
$\leq$ $C\Delta^{(1-\alpha)p}.$
Using the same technique as in (3.3), along with the Hölder inequality, the Fubini theorem and Lemmas 2.5, 5.4, we get
$J_{2}\leq C\Delta^{pH-1}\int_{t_{k}}^{t_{k+1}}\mathbb{E}(1+\mathbb{W}_{\theta}(\overline{\mu}_{s}^{Y,N},\delta_{0}))^{p}{\rm d}s\leq C\Delta^{pH}.$
Recalling that $\alpha\in(0,1/2]$, we obtian
$\mathbb{E}\bigg{(}\sup_{t_{k}\leq t\leq t_{k+1}}|\widetilde{Y}_{\Delta}^{i,N}(t)-\widetilde{Y}_{\Delta}^{i,N}(t_{k})|^{p}\bigg{)}\leq C\Delta^{[(1-\alpha)\wedge H]p}.$
Denoting by $\widetilde{D}(t,t_{k}):=D(Y_{\Delta}^{i,N}(t-\tau))-D(Y_{\Delta}^{i,N}(t_{k}-\tau))$, and $\widetilde{b}(t,t_{k}):=b_{\Delta}(Y_{\Delta}^{i,N}(t),Y_{\Delta}^{i,N}(t-\tau),\mu^{Y,N}(t))-b_{\Delta}(Y_{\Delta}^{i,N}(t_{k}),Y_{\Delta}^{i,N}(t_{k}-\tau),\mu^{Y,N}(t_{k}))$, together with Assumption 2.3, we get
(5.13) $|\widetilde{Y}_{\Delta}^{i,N}(t)-\widetilde{Y}_{\Delta}^{i,N}(t_{k})|^{p}$
$\geq$ $2^{1-p}|Y_{\Delta}^{i,N}(t)-Y_{\Delta}^{i,N}(t_{k})-\widetilde{D}(t,t_{k})|^{p}-\theta^{p}\Delta^{p}|\widetilde{b}(t,t_{k})|^{p}$
$\geq$ $2^{2-2p}|Y_{\Delta}^{i,N}(t)-Y_{\Delta}^{i,N}(t_{k})|^{p}-2^{1-p}\lambda^{p}|Y_{\Delta}^{i,N}(t-\tau)-Y_{\Delta}^{i,N}(t_{k}-\tau)|^{p}-\theta^{p}\Delta^{p}|\widetilde{b}(t,t_{k})|^{p}.$
For $0\leq t\leq t_{1}=\Delta$, then $t-\tau<t_{1}-\tau<0$, we derive from (5.13) and Lemma 5.4,
$\mathbb{E}\bigg{(}\sup_{0\leq t\leq t_{1}}|Y_{\Delta}^{i,N}(t)-Y_{\Delta}^{i,N}(t_{0})|^{p}\bigg{)}\leq C\mathbb{E}\bigg{(}\sup_{0\leq t\leq t_{1}}|\widetilde{Y}_{\Delta}^{i,N}(t)-\widetilde{Y}_{\Delta}^{i,N}(t_{0})|^{p}\bigg{)}+C\Delta^{(1-\alpha)p}\leq C\Delta^{[(1-\alpha)\wedge H]p}.$
Similarly, for $t_{1}\leq t<t_{2}$, we obtain
$\mathbb{E}\bigg{(}\sup_{t_{1}\leq t\leq t_{2}}|Y_{\Delta}^{i,N}(t)-Y_{\Delta}^{i,N}(t_{1})|^{p}\bigg{)}$
$\leq$ $C\mathbb{E}\bigg{(}\sup_{t_{1}\leq t\leq t_{2}}|\widetilde{Y}_{\Delta}^{i,N}(t)-\widetilde{Y}_{\Delta}^{i,N}(t_{1})|^{p}\bigg{)}+C\mathbb{E}\bigg{(}\sup_{0\leq t\leq(t_{2-m}\vee 0)}|\widetilde{Y}_{\Delta}^{i,N}(t)-\widetilde{Y}_{\Delta}^{i,N}(t_{1-m})|^{p}\bigg{)}+C\Delta^{(1-\alpha)p}$
$\leq$ $C\Delta^{[(1-\alpha)\wedge H]p}.$
By induction,
$\mathbb{E}\bigg{(}\sup_{t_{k}\leq t\leq t_{k+1}}|Y_{\Delta}^{i,N}(t)-Y_{\Delta}^{i,N}(t_{1})|^{p}\bigg{)}\leq$ $C\mathbb{E}\bigg{(}\sup_{t_{k}\leq t\leq t_{k+1}}|\widetilde{Y}_{\Delta}^{i,N}(t)-\widetilde{Y}_{\Delta}^{i,N}(t_{k})|^{p}\bigg{)}+C\Delta^{(1-\alpha)p}$
$\leq$ $C\Delta^{[(1-\alpha)\wedge H]p}.$
we get the desired assertion. ∎
Lemma 5.6.
Let Assumptions 2.2-2.5 and Lemmas 5.1, 5.2 hold, then it holds that for $p>1/H$ and $\overline{p}>2(l+1)p$,
$\mathbb{E}\bigg{(}\sup_{0\leq t\leq T}|X_{t}^{i,N}-Y_{\Delta}^{i,N}(t)|^{p}\bigg{)}\leq C\Delta^{\alpha p},$
where $C$ is a positive constant independent of $\Delta$.
Proof.
For $t\in[0,T]$, denote $I(t)=\widetilde{Y}_{\Delta}^{i,N}(t)-X_{t}^{i,N}+D(X_{t-\tau}^{i,N})$, then
$I(t)=$ $I(0)+\int_{0}^{t}(b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})){\rm d}s$
$+\int_{0}^{t}(\sigma(\overline{\mu}^{Y,N}(s))-\sigma(\mu_{s}^{X,N})){\rm d}B_{s}^{H,i},$
where $I_{0}=-\theta b_{\Delta}(\xi(0),\xi(-\tau),\delta_{0})\Delta$. The Itô formula yields,
$|I(t)|^{p}\leq$ $|I(0)|^{p}$
$+p\int_{0}^{t}|I(s)|^{p-2}\langle I(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})\rangle{\rm d}s$
$+HT^{2H-1}p(p-1)\int_{0}^{t}|I(s)|^{p-2}|\sigma(\overline{\mu}^{Y,N}(s))-\sigma(\mu_{s}^{X,N})|^{2}{\rm d}s$
$+p\int_{0}^{t}|I(s)|^{p-2}\langle I(s),\sigma(\overline{\mu}^{Y,N}(s))-\sigma(\mu_{s}^{X,N})\rangle{\rm d}B_{s}^{H,i}$
$\leq$ $|I(0)|^{p}+\sum_{i=1}^{6}H_{i}(t),$
where
$H_{1}(t)=$ $p\int_{0}^{t}|I(s)|^{p-2}\langle I(s),b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))$
$-b(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))\rangle{\rm d}s$
$H_{2}(t)=$ $p\int_{0}^{t}|I(s)|^{p-2}\langle I(s),b(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))$
$-b(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))\rangle{\rm d}s$
$H_{3}(t)=$ $p\int_{0}^{t}|I(s)|^{p-2}\langle I(s),b(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))$
$-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})\rangle{\rm d}s,$
$H_{4}(t)=$ $HT^{2H-1}p(p-1)\int_{0}^{t}|I(s)|^{p-2}|\sigma(\overline{\mu}^{Y,N}(s))-\sigma(\mu^{Y,N}(s))|^{2}{\rm d}s$
$H_{5}(t)=$ $HT^{2H-1}p(p-1)\int_{0}^{t}|I(s)|^{p-2}|\sigma(\mu^{Y,N}(s))-\sigma(\mu_{s}^{X,N})|^{2}{\rm d}s$
$H_{6}(t)=$ $p\int_{0}^{t}|I(s)|^{p-2}\langle I(s),\sigma(\overline{\mu}^{Y,N}(s))-\sigma(\mu_{s}^{X,N})\rangle{\rm d}B_{s}^{H,i}.$
By Assumption 2.3, Lemmas 5.1, 5.2, 5.4, and the Young’s inequality, the Hölder inequality, we get
$\mathbb{E}\bigg{(}\sup_{0\leq u\leq t}H_{1}(u)\bigg{)}$
$\leq$ $C\mathbb{E}\int_{0}^{t}|I(s)|^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}|b_{\Delta}(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))$
$-b(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))|^{p}{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}\bigg{(}|Y_{\Delta}^{i,N}(s)-X_{s}^{i,N}|^{p}+|Y_{\Delta}^{i,N}(s-\tau)-X_{s-\tau}^{i,N}|^{p}$
$+\theta^{p}\Delta^{p}|b_{\Delta}(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))|^{p}\bigg{)}{\rm d}s$
$+C\Delta^{\alpha p}\mathbb{E}\int_{0}^{t}(1+|\overline{Y}_{\Delta}^{i,N}(s)|^{2(l+1)p}+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|^{2(l+1)p}+(\mathbb{W}_{q}(\overline{\mu}^{Y,N},\delta_{0}))^{2p}){\rm d}s$
$\leq$ $C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(s)-X_{s}^{i,N}|^{p}+C\Delta^{(1-\alpha)p}+C\Delta^{\alpha p},$
where for the last term we used the fact that $\overline{p}>2(l+1)p$. By the elementary inequality, the Young’s inequality, the Hölder inequality, Assumptions 2.3, 2.4 and Lemmas 5.1, 5.4, 5.5, we get
$\mathbb{E}\bigg{(}\sup_{0\leq u\leq t}H_{2}(u)\bigg{)}$
$\leq$ $C\mathbb{E}\int_{0}^{t}|I(s)|^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}|b(\overline{Y}_{\Delta}^{i,N}(s),\overline{Y}_{\Delta}^{i,N}(s-\tau),\overline{\mu}^{Y,N}(s))-b(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))|^{p}{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}\bigg{(}|Y_{\Delta}^{i,N}(s)-X_{s}^{i,N}|^{p}+|Y_{\Delta}^{i,N}(s-\tau)-X_{s-\tau}^{i,N}|^{p}+\theta^{p}\Delta^{p}|b_{\Delta}(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))|\bigg{)}{\rm d}s$
$+C\mathbb{E}\int_{0}^{t}\bigg{[}(1+|\overline{Y}_{\Delta}^{i,N}(s)|^{l}+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|^{l}+|Y_{\Delta}^{i,N}(s)|^{l}+|Y_{\Delta}^{i,N}(s-\tau)|^{l})^{p}$
×$\times(|\overline{Y}_{\Delta}^{i,N}(s)-Y_{\Delta}^{i,N}(s)|+|\overline{Y}_{\Delta}^{i,N}(s-\tau)-Y_{\Delta}^{i,N}(s-\tau)|+(\mathbb{W}_{q}(\overline{\mu}^{Y,N}(s),\mu^{Y,N})(s)))^{p}\bigg{]}{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}\bigg{(}|Y_{\Delta}^{i,N}(s)-X_{s}^{i,N}|^{p}+|Y_{\Delta}^{i,N}(s-\tau)-X_{s-\tau}^{i,N}|^{p}\bigg{)}{\rm d}s+C\Delta^{(1-\alpha)p}$
$+C\int_{0}^{t}[\mathbb{E}(1+|\overline{Y}_{\Delta}^{i,N}(s)|^{l}+|\overline{Y}_{\Delta}^{i,N}(s-\tau)|^{l}+|Y_{\Delta}^{i,N}(s)|^{l}+|Y_{\Delta}^{i,N}(s-\tau)|^{l})^{2p}]^{1/2}$
×$\times[\mathbb{E}(|\overline{Y}_{\Delta}^{i,N}(s)-Y_{\Delta}^{i,N}(s)|+|\overline{Y}_{\Delta}^{i,N}(s-\tau)-Y_{\Delta}^{i,N}(s-\tau)|)^{2p}]^{1/2}{\rm d}s$
$+C\int_{0}^{t}\mathbb{E}(|\overline{Y}_{\Delta}^{i,N}(s)-Y_{\Delta}^{i,N}(s)|)^{p}{\rm d}s$
$\leq$ $C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}{\rm d}s+C\Delta^{(1-\alpha)p}+C\Delta^{[(1-\alpha)\wedge H]p}.$
By Assumption 2.4 and Lemmas 5.1, 5.4, we get
$\mathbb{E}\bigg{(}\sup_{0\leq u\leq t}H_{3}(u)\bigg{)}$
$\leq$ $C\mathbb{E}\int_{0}^{t}|I(s)|^{p-2}[|Y_{\Delta}^{i,N}(s)-X_{s}^{i,N}|^{2}+|Y_{\Delta}^{i,N}(s-\tau)-X_{s-\tau}^{i,N}|^{2}+(\mathbb{W}_{q}(\mu^{Y,N}(s),\mu^{X,N}(s)))^{2}]{\rm d}s$
$+C\mathbb{E}\int_{0}^{t}|I(s)|^{p-2}|\theta b_{\Delta}(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))\Delta||b(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))$
$-b(X_{s}^{i,N},X_{s-\tau}^{i,N},\mu_{s}^{X,N})|{\rm d}s$
$\leq$ $C\mathbb{E}\int_{0}^{t}[|Y_{\Delta}^{i,N}(s)-X_{s}^{i,N}|^{p}+\theta^{p}\Delta^{p}|b_{\Delta}(Y_{\Delta}^{i,N}(s),Y_{\Delta}^{i,N}(s-\tau),\mu^{Y,N}(s))|^{p}]{\rm d}s$
$+C\Delta^{1-\alpha}\mathbb{E}\int_{0}^{t}|I(s)|^{p-2}[(1+|Y_{\Delta}^{i,N}(s)|^{l}+|Y_{\Delta}^{i,N}(s-\tau)|^{l}+|X_{s}^{i,N}|^{l}+|X_{s-\tau}^{i,N}|^{l})$
×$\times(|Y_{\Delta}^{i,N}(s)-X_{s}^{i,N}|+|Y_{\Delta}^{i,N}(s-\tau)-X_{s-\tau}^{i,N}|)+\mathbb{W}_{q}(\mu^{Y,N}(s),\mu_{s}^{X,N})]{\rm d}s$
$\leq$ $C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}{\rm d}s+C\Delta^{(1-\alpha)(1+p)}.$
By the Young’s inequality, Assumption 2.5, we get
$\mathbb{E}\bigg{(}\sup_{0\leq u\leq t}H_{4}(u)\bigg{)}+\mathbb{E}\bigg{(}\sup_{0\leq u\leq t}H_{5}(u)\bigg{)}$
$\leq$ $C\mathbb{E}\int_{0}^{t}|I(s)|^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}(\mathbb{W}_{q}(\overline{\mu}^{Y,N}(s)),\mu^{Y,N}(s))^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}(\mathbb{W}_{q}(\mu^{Y,N}(s)),\mu_{s}^{X,N})^{p}{\rm d}s$
$\leq$ $C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}{\rm d}s+C\Delta^{(1-\alpha)p}+C\Delta^{[(1-\alpha)\wedge H]p}.$
In the way in (3.3), by the Young’s inequality, we get
$\mathbb{E}\bigg{(}\sup_{0\leq u\leq t}|H_{6}(u)|^{p}\bigg{)}$
$\leq$ $C\mathbb{E}\int_{0}^{t}|I(s)|^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}(\mathbb{W}_{q}(\overline{\mu}^{Y,N}(s)),\mu^{Y,N}(s))^{p}{\rm d}s+C\mathbb{E}\int_{0}^{t}(\mathbb{W}_{q}(\mu^{Y,N}(s)),\mu_{s}^{X,N})^{p}{\rm d}s$
$\leq$ $C\int_{0}^{t}\sup_{0\leq u\leq s}\mathbb{E}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}{\rm d}s+C\Delta^{(1-\alpha)p}+C\Delta^{[(1-\alpha)\wedge H]p}.$
Combining $H_{1}-H_{6}$, we get
$\mathbb{E}\bigg{(}\sup_{0\leq u\leq t}|I(u)|^{p}\bigg{)}\leq C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{0\leq u\leq s}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}\bigg{)}{\rm d}s+C\Delta^{\alpha p}.$
Recalling the definition of $I(t)$ and by Assumption 2.3, we may compute
$|I(t)|^{p}\geq$ $2^{1-p}|Y_{\Delta}^{i,N}(t)-X_{t}^{i,N}-D(Y_{\Delta}^{i,N}(t-\tau))+D(X_{t-\tau}^{i,N})|^{p}$
$-|\theta b_{\Delta}(Y_{\Delta}^{i,N}(t),Y_{\Delta}^{i,N}(t-\tau),\mu^{Y,N}(t))\Delta|^{p}$
$\geq$ $2^{1-p}[2^{1-p}|Y_{\Delta}^{i,N}(t)-X_{t}^{i,N}|^{p}-|D(Y_{\Delta}^{i,N}(t-\tau))-D(X_{t-\tau}^{i,N})|^{p}]$
$-|\theta b_{\Delta}(Y_{\Delta}^{i,N}(t),Y_{\Delta}^{i,N}(t-\tau),\mu^{Y,N}(t))\Delta|^{p}$
$\geq$ $2^{2-2p}|Y_{\Delta}^{i,N}(t)-X_{t}^{i,N}|^{p}-2^{1-p}\lambda^{p}|Y_{\Delta}^{i,N}(t-\tau)-X_{t-\tau}^{i,N}|^{p}$
$-|\theta b_{\Delta}(Y_{\Delta}^{i,N}(t),Y_{\Delta}^{i,N}(t-\tau),\mu^{Y,N}(t))\Delta|^{p}.$
By Lemma 5.1, 5.4, we derive
$\mathbb{E}\bigg{(}\sup_{0\leq u\leq t}|X_{u}^{i,N}-Y_{\Delta}^{i,N}(u)|^{p}\bigg{)}\leq$ $\mathbb{E}\bigg{(}\sup_{0\leq u\leq t}|I(u)|^{p}\bigg{)}+C\Delta^{(1-\alpha)p}$
$\leq$ $C\int_{0}^{t}\mathbb{E}\bigg{(}\sup_{0\leq u\leq s}|Y_{\Delta}^{i,N}(u)-X_{u}^{i,N}|^{p}\bigg{)}{\rm d}s+C\Delta^{\alpha p}.$
Applying the Gronwall inequality, we can derive the assertion. ∎
Theorem 5.7.
Let Assumptions 2.2-2.5 hold. For $p>1/H$ and $\overline{p}\geq 2(l+1)p$, it holds that
$\mathbb{E}\bigg{(}\sup_{t\in[0,T]}|X_{t}^{i}-Y_{t}^{i.N}|^{p}\bigg{)}\leq C\times\left\{\begin{array}[]{ll}N^{-\frac{1}{2}}+\Delta^{\alpha p},\ \ p>\frac{d}{2},\\ N^{-\frac{1}{2}}\log(1+N)+\Delta^{\alpha p},\ p=\frac{d}{2},\\ N^{-\frac{p}{d}}+\Delta^{\alpha p},\ p\in[2,\frac{d}{2}),\end{array}\right.$
where $C$ is a constant independent of $N$ and $\Delta$.
Proof.
By Lemmas 4.3 and 5.6, the result is obvious. ∎